### v2

# Theoretical Guarantees for the Subspace-Constrained Tyler’s Estimator

Abstract
This work analyzes the subspace-constrained Tyler’s estimator (STE) [12] designed for recovering a low-dimensional subspace within a dataset that may be highly corrupted with outliers. It assumes a weak inlier-outlier model and allows the fraction of inliers to be smaller than a fraction that leads to computational hardness of the robust subspace recovery problem. It shows that in this setting, if the initialization of STE, which is an iterative algorithm, satisfies a certain condition, then STE can effectively recover the underlying subspace. It further shows that under the generalized haystack model, STE initialized by the Tyler’s M-estimator (TME), can recover the subspace when the fraction of iniliers is too small for TME to handle.
1 Introduction
The subspace-constrained Tyler’s estimator (STE) is a recent method for robust subspace recovery (RSR), which was recently suggested in [12]. In particular, [12] states two theorems for STE. While [12] addresses the computer vision community, we believe that the proofs of these theorems are of independent interest to some mathematicians and thus include them in this preprint, which will eventually evolve into a complete manuscript.
The rest of this introduction reviews STE and its stated theory: §1.1 provides very basic notations, §1.2 reviews the STE algorithm of [12] and §1.3 reviews the statement of its theory, with their required definitions and notation. Since this is a review of [12] we repeat some of its text at this current version, which supplements [12]. In the rest of the paper, we clarify a technical assumption of the theory in §2 and prove the stated theorems in §3.
1.1 Notation
We use bold upper and lower case letters for matrices and column vectors, respectively. Let ${\mathbf{I}}_{k}$ denote the identity matrix in $\mathbb{R}^{k\times k}$, where if $k$ is obvious we just write ${\mathbf{I}}$. For a matrix ${\mathbf{A}}$, we denote by $\operatorname{tr}({\mathbf{A}})$ and $\operatorname{Im}({\mathbf{A}})$ the trace and image (i.e., column space) of ${\mathbf{A}}$. We denote by $S_{+}(D)$ and $S_{++}(D)$ the sets of positive semidefinite and definite matrices in $\mathbb{R}^{D\times D}$, respectively. We denote by $O(D,d)$ the set of semi-orthogonal ${D\times d}$ matrices, i.e., ${\mathbf{U}}\in O(D,d)$ if and only if ${\mathbf{U}}\in\mathbb{R}^{D\times d}$ and ${\mathbf{U}}^{\top}{\mathbf{U}}={\mathbf{I}}_{d}$. We refer to linear $d$-dimensional subspaces as $d$-subspaces. For a $d$-subspace $L$, we denote by ${\mathbf{P}}_{L}$ the $D\times D$ matrix representing the orthogonal projector onto $L$. We also arbitrarily fix ${\mathbf{U}}_{L}$ in $O(D,d)$ such that ${\mathbf{U}}_{L}{\mathbf{U}}_{L}^{\top}={\mathbf{P}}_{L}$ (such ${\mathbf{U}}_{L}$ is determined up to right multiplication by an orthogonal matrix in $O(d,d)$). Throughout the paper, ${\mathcal{X}}=\{{\bm{x}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{D}$ is assumed to be a given centered dataset, that is, $\sum_{i=1}^{N}{\bm{x}}_{i}={\mathbf{0}}$.
1.2 Review of the STE Algorithm
The STE algorithm [12] aims to use more cleverly the $d$-subspace information within the TME framework, instead of first estimating the full covariance. At iteration $k$ it follows a similar step to that of TME:
${\mathbf{Z}}^{(k)}:=\sum_{i=1}^{N}{\bm{x}}_{i}{\bm{x}}_{i}^{\top}/({\bm{x}}_{i}^{\top}({\mathbf{\Sigma}}^{(k-1)})^{-1}{\bm{x}}_{i}).$
It computes the eigenvalues $\{\sigma_{i}\}_{i=1}^{D}$ of ${\mathbf{Z}}^{(k)}$ and replaces each of the bottom $(D-d)$ of them with $\gamma\sigma_{d+1,D}$, where
$\sigma_{d+1,D}:=\frac{1}{D-d}\sum_{i=d+1}^{D}\sigma_{i}.$ (1)
It also computes the eigenvectors of ${\mathbf{Z}}^{(k)}$ and forms the matrix ${\mathbf{\Sigma}}^{(k)}$ with the same eigenvectors as those of ${\mathbf{Z}}^{(k)}$ and the modified eigenvalues, scaled to have trace 1. It iteratively repeats this procedure until the two estimators are sufficiently close. Algorithm 1 summarizes this procedure. By observing this algorithm one can note that STE is invariant to scaling of the data.
Algorithm 1 STE: Subspace-Constrained Tyler’s Estimator
1:  Input: ${\mathbf{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{N}]\in{\mathbb{R}}^{D\times N}$: centered data matrix, $d$: subspace dimension, $K$: maximum number of iterations, $\tau,\gamma$: parameters.
2:  Output: $L$: $d$-subspace in ${\mathbb{R}}^{D}$
3:  ${\mathbf{\Sigma}}^{(0)}={\mathbf{I}}_{D}/D$
4:  for $k=1,2,\ldots$ do
5:     ${\mathbf{Z}}^{(k)}\leftarrow\sum_{i=1}^{N}{\bm{x}}_{i}{\bm{x}}_{i}^{\top}/({\bm{x}}_{i}^{\top}({\mathbf{\Sigma}}^{(k-1)})^{-1}{\bm{x}}_{i})$
6:     $[{\mathbf{U}}^{(k)},{\mathbf{S}}^{(k)},{\mathbf{U}}^{(k)}]\leftarrow\text{EVD}({\mathbf{Z}}^{(k)})$
7:     $\sigma_{i}\leftarrow[{\mathbf{S}}^{(k)}]_{ii}$ and $\sigma_{d+1,D}\leftarrow\sum_{i=d+1}^{D}{\sigma_{i}}/{(D-d)}$
8:     $\widetilde{{\mathbf{S}}}^{(k)}\leftarrow\operatorname{diag}(\sigma_{1},\ldots,\sigma_{d},\gamma\cdot\sigma_{d+1,D},\ldots,\gamma\cdot\sigma_{d+1,D})$,
9:     ${\mathbf{\Sigma}}^{(k)}\leftarrow{\mathbf{U}}^{(k)}\widetilde{{\mathbf{S}}}^{(k)}({\mathbf{U}}^{(k)})^{\top}/\operatorname{tr}\big{(}{\mathbf{U}}^{(k)}\widetilde{{\mathbf{S}}}^{(k)}({\mathbf{U}}^{(k)})^{\top}\big{)}$
10:     Stop if $k\geq K$ or $\|{\mathbf{\Sigma}}^{(k)}-{\mathbf{\Sigma}}^{(k-1)}\|_{F}<\tau$
11:  end for
12:  $L$ = Span of the first $d$ columns of ${\mathbf{U}}^{(k)}$
1.3 Review of the Stated Theory
Generic Model: The generic STE theory assumes a noiseless inliers-outliers RSR model. Let $L_{*}$ denote the underlying $d$-subspace in $\mathbb{R}^{D}$, ${\mathcal{X}}_{in}={\mathcal{X}}\cap L_{*}$ and ${\mathcal{X}}_{out}={\mathcal{X}}\setminus{\mathcal{X}}_{in}$ be the set of inliers and outliers, respectively, and $n_{1}=|{\mathcal{X}}_{in}|$ and $n_{0}=|{\mathcal{X}}_{out}|$ be the number of inliers and outliers. The first assumption of the generic STE theory is
Assumption 1: Any $k$-subspace of $L_{*}$, $1\leq k\leq d$, contains at most $n_{1}k/d$ points.
The next assumption uses the following notion of dimension-scaled signal-to-noise ratio. In RSR, the ratio of inliers per outliers, $n_{1}/n_{0}$, is often referred to as the SNR (signal-to-noise ratio) [8, 5, 7]. The dimension-scaled SNR (DS-SNR) [12] is obtained when scaling $n_{1}$ and $n_{0}$ by their respective dimensions (of $L_{*}$ and $L_{*}^{\perp}$):
[…]
Zhang [13] showed that exact recovery by TME is guaranteed whenever DS-SNR$>1$ (assuming general position assumptions on the inliers and outliers). In fact, our theoretical estimates show that it is okay to assume instead DS-SNR1absent1$\geq$ 1≥ 1. Furthermore, Hardt and Moitra [2] showed that when considering general datasets with general position assumptions on the inliers and outliers, the RSR problem is SSE hard if the DS-SNR is lower than $1$. The second assumption of the generic STE theory is
[…]
The last assumption for the generic theory of STE requires a sufficiently good initialization for STE, but also implicitly involves additional hidden assumptions on the inliers and outliers. To formulate the new assumption we define below some some basic condition numbers for good initialization (which are more complicated than the one for initialization by PCA suggested by [8] and [7]) and also quantities similar to the ones used to guarantee landscape stability in the theory of RSR [6, 14, 8, 5].
Definitions required for Assumption 3: Recall that ${\mathbf{\Sigma}}^{(0)}$ denotes the initial value in Algorithm 1, and denote
${\mathbf{\Sigma}}^{(0)}_{L_{1},L_{2}}={\mathbf{U}}_{L_{1}}^{\top}{\mathbf{\Sigma}}^{(0)}{\mathbf{U}}_{L_{2}}.$
We define the following condition number
$\kappa_{1}=\frac{\sigma_{d}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*},L_{*}}-{\mathbf{\Sigma}}^{(0)}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{(0)\,-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}}\Big{)}}{\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}}.$
To get a better intuition to this primary quantity of Assumption 3, we first express the initial estimator ${\mathbf{\Sigma}}^{(0)}$, using basis vectors for $L_{*}$ and $L_{*}^{\perp}$, as a $2\times 2$ block matrix
$\begin{pmatrix}{\mathbf{\Sigma}}^{(0)}_{L_{*},L_{*}}&{\mathbf{\Sigma}}^{(0)}_{L_{*},L_{*}^{\perp}}\\ {\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}.$
Defining ${\mathbf{\Sigma}}^{\prime}={\mathbf{\Sigma}}^{(0)}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{(0)\,-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}}$, we decompose this block matrix as
$\begin{pmatrix}{\mathbf{\Sigma}}^{\prime}&{\mathbf{\Sigma}}^{(0)}_{L_{*},L_{*}^{\perp}}\\ {\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}},\end{pmatrix}+\begin{pmatrix}{\mathbf{\Sigma}}^{(0)}_{L_{*},L_{*}}-{\mathbf{\Sigma}}^{\prime}&0\\ 0&0\end{pmatrix}.$
We note that the numerator of $\kappa_{1}$ is the $d$-th eigenvalue of the second matrix in the above sum. We show in Lemma 1(a) that this eigenvalue is positive if ${\mathbf{\Sigma}}^{(0)}$ is positive definite, which can be easily enforced. The condition number is thus the ratio between the smallest positive eigenvalue of the second matrix of the sum and the largest eigenvalue of the component of the first matrix associated with $L_{*}^{\perp}$. Therefore, $\kappa_{1}$ expresses a ratio between a quantifier of a $d$-dimensional component of ${\mathbf{\Sigma}}^{(0)}$, associated with $L_{*}$, and a quantifier of the projection onto $L_{*}^{\perp}$ of a full rank component of ${\mathbf{\Sigma}}^{(0)}$.
We also define ${\mathbf{\Sigma}}_{in,*}$ as the TME solution to the set of the projected inliers $\{{\mathbf{U}}_{L^{*}}{\bm{x}}\mid{\bm{x}}\in{\mathcal{X}}_{in}\}\subset{\mathbb{R}}^{d}$ and the following two condition numbers
$\kappa_{2}=\frac{\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}}{\sigma_{D}({\mathbf{\Sigma}}^{(0)})}\ \text{ and }\ \kappa_{in}=\frac{\sigma_{1}({\mathbf{\Sigma}}_{in,*})}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}.$
We note that $\kappa_{in}$ is analogous to the one in (25) of [7], where we replace the sample covariance by the TME estimator. An analog to the alignment of outliers statistic [6, 8] for STE is
${\mathcal{A}}=\Big{\|}\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}.$
An analog to the stability statistic [6, 8] for STE is
${\mathcal{S}}=\sigma_{d+1,D}\Big{(}\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}\Big{)},$
where $\sigma_{d+1,D}({\mathbf{X}})$ was defined in (1).
[…]
$\kappa_{1}\geq C\,\frac{d\,\kappa_{in}\,{\mathcal{A}}}{n_{1}}\left(\kappa_{in}+\frac{{\mathcal{A}}}{\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}}+\frac{\kappa_{2}{\mathcal{A}}}{\gamma{\mathcal{S}}}(1+\kappa_{in})\right).$ (3)
The exact technical requirement on $C$ is specified in our proof. In general, the larger the RHS of (3), the more restricted the choice of ${\mathbf{\Sigma}}^{(0)}$ is. In particular, when $\kappa_{1}=\infty$, the definition of $\kappa_{1}$ implies that $\operatorname{Im}({\mathbf{\Sigma}}^{(0)})=L_{*}$, so the subspace is already recovered by the initial estimate. Therefore, reducing the lower bound of $\kappa_{1}$ may allow some flexibility, so a marginally suboptimal initialization could still work out. In §2 we show that under the asymptotic generalized haystack model, Assumption 3 can be interpreted as an upper bound on the largest principal angle between the initial and ground truth subspaces.
Generic Theory: The next theorem suggests that under assumptions 1-3, STE nicely converges to an estimator that recovers $L_{*}$. The main significance of this theory is that its assumptions can allow DS-SNR lower than 1 for special instances of datasets (for which the assumptions hold), unlike the general recovery theories of [2] and [13]. The main significance of this theory is that its assumptions can allow DS-SNR lower than 1 for special instances of datasets (for which the assumptions hold), unlike the general recovery theories of [13] and [2].
Theorem 1.
Under assumptions 1-3, the sequence ${\mathbf{\Sigma}}^{(k)}$ generated by STE converges to ${\mathbf{U}}_{L_{*}}{\mathbf{\Sigma}}_{in,*}{\mathbf{U}}_{L_{*}}^{\top}$, the TME solution for the set of inliers ${\mathcal{X}}_{in}$. In addition, let $L^{(k)}$ be the subspace spanned by the top $d$ eigenvectors of ${\mathbf{\Sigma}}^{(k)}$, then the angle between $L^{(k)}$ and $L_{*}$, $\angle(L^{(k)},L_{*})=\cos^{-1}(\|{\mathbf{U}}_{L^{(k)}}^{\top}{\mathbf{U}}_{L_{*}}\|)$, converges $r$-linearly to zero.
We discuss insights of this theory on choices of the algorithms and further verify the above stated advantage of STE over TME assuming a common probabilistic model.
Possible ways of Initialization: If one expects an initial estimated subspace $\hat{L}$ to have a sufficiently small angle $\theta$ with $L_{*}$, where $\theta=\angle(\hat{L},L_{*})$, then for ${\mathbf{\Sigma}}^{(0)}:=\Pi_{\hat{L}}+\epsilon{\mathbf{I}}$ it can be shown that $\kappa_{1}>O(1/(\epsilon+\theta))$ and $\kappa_{2}<O(1+\frac{\theta}{\epsilon})$. Thus one may use a trusted RSR method, e.g., FMS. The choice ${\mathbf{\Sigma}}^{(0)}={\mathbf{I}}$ (or a scaled version of it) corresponds to $\hat{L}$ being the PCA subspace (obtain at iteration 1). Also, using the TME solution for ${\mathbf{\Sigma}}^{(0)}$ corresponds to using the TME subspace as $\hat{L}$.
Theory under a probabilistic model: We show that under a common probabilistic model, the assumptions of Theorem 1, where ${\mathbf{\Sigma}}^{(0)}$ is obtained by TME, hold. Moreover, we show that STE (initialized by TME) can recover the correct subspace in situations with DS-SNR$<1$, whereas TME cannot recover the underlying subspace in such cases. We follow [8] and study the Generalized Haystack Model, though for simplicity, we assume Gaussian instead of sub-Gaussian distributions and an asymptotic setting. We assume $n_{1}$ inliers i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(in)}/d)$, where ${\mathbf{\Sigma}}^{(in)}\in S_{+}(D)$ and $L_{*}=\operatorname{Im}({\mathbf{\Sigma}}^{(in)})$ (so ${\mathbf{\Sigma}}^{(in)}$ has $d$ nonzero eigenvalues), and $n_{0}$ outliers are i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(out)}/D)$, where ${\mathbf{\Sigma}}^{(out)}/D\in S_{++}(D)$. We define the following condition numbers of inliers (in $L_{*}$) and outliers:
$\kappa_{in}=\frac{\sigma_{1}({\mathbf{\Sigma}}^{(in)})}{\sigma_{d}({\mathbf{\Sigma}}^{(in)})}\,\,\ \ \text{ and }\ \kappa_{out}=\frac{\sigma_{1}({\mathbf{\Sigma}}^{(out)})}{\sigma_{D}({\mathbf{\Sigma}}^{(out)})}\,.$
Clearly, Assumption 1 holds under this model, and Assumption 2 constrains some of its parameters. Our next theorem shows that Assumption 3 holds under this model when the initial estimate ${\mathbf{\Sigma}}^{(out)}$ for STE is obtained by TME. It also shows that in this case STE can solve the RSR problem even when DS-SNR$<1$, unlike TME. For simplicity, we formulate the theory for the asymptotic case, where $N\rightarrow\infty$ and the theorem holds almost surely. It is possible to formulate it for a very large $N$ with high probability, but it requires stating complicated constants depending on various parameters.
Theorem 2.
[…]
There are three different regimes that the theorem covers. When DS-SNR1absent1$\geq$ 1≥ 1, both TME+STE (i.e., STE initialized by TME) and TME solve the RSR problem. When $\eta\leq$DS-SNR$<1$, TME+STE solves the RSR problem and TME generally fails. When $\gamma\leq$DS-SNR$<\eta$, TME+STE might also fail, but STE with extremely good initialization (that satisfies Assumption 3) can still solve the problem.
To get a basic idea of the dependence of $\eta$ on its parameters, we remark that $\eta\rightarrow 1$ if either $c_{0}\rightarrow 0$, $\kappa_{in}\rightarrow\infty$, $\kappa_{out}\rightarrow\infty$ or $\mu\rightarrow 0$, where the parameter $\mu$ is somewhat artificial and might be removed with a tighter proof. Therefore, successful performance of TME+STE requires a DS-SNR that is close to $1$ when $\gamma$ is close to either 00 or $\eta$ (so that $c_{0}$ is very small) or when either the inlier or outlier distribution is highly non-symmetric, that is, either $\kappa_{in}$ or $\kappa_{out}$ is large.
2 Clarification of Assumption 3 under the Asymptotic Generalized Haystack Model
We try to comprehend the technical expression in the RHS of (3), which appear in Assumption 3, under the Asymptotic Generalized Haystack.
We follow [8] and study the Generalized Haystack Model, though for simplicity, we assume Gaussian instead of sub-Gaussian distributions and we also assume an asymptotic setting with sufficiently large numbers of inliers and outliers, though the overall fraction of inliers can be small. We assume that $n_{1}$ inliers are i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(in)}/d)$, where ${\mathbf{\Sigma}}^{(in)}\in S_{+}(D)$ and $L_{*}=\operatorname{Im}({\mathbf{\Sigma}}^{(in)})$ (so ${\mathbf{\Sigma}}^{(in)}$ has $d$ nonzero eigenvalues), and $n_{0}$ percentage of the outliers are i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(out)}/D)$, where ${\mathbf{\Sigma}}^{(out)}\in S_{++}(D)$.
Let $\mu_{out}$ denote the projection of $N(0,{\mathbf{\Sigma}}^{(out)}/D)$ on the unit sphere, that is, it is the pushforward measure of $N(0,{\mathbf{\Sigma}}^{(out)}/D)$ induced by the projection to the unit sphere in ${\mathbb{R}}^{D}$. We note that
$\lim_{n_{0}\rightarrow\infty}\frac{{\mathcal{A}}}{n_{0}}=\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{\Sigma}}^{(out)}/D)}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}=\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim\mu_{out}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}.$ (4)
Indeed, the first equality follows from the definition of ${\mathcal{A}}$ and the law of large numbers and the second one follows from the invariance to scaling of ${\mathbf{x}\mathbf{x}^{\top}}/{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}$.
Let $\mu_{0}$ denote the uniform distribution on the unit sphere in ${\mathbb{R}}^{D}$. Let $f_{\mu_{out}}$ and $f_{\mu_{0}}$ denote the probability density functions of $\mu_{out}$ and $\mu_{0}$. Furthermore, let $C_{\mu_{out},u},C_{\mu_{out},l}>0$ denote the optimal constants such that
$C_{\mu_{out},l}f_{\mu_{0}}\leq f_{\mu_{out}}\leq C_{\mu_{out},u}f_{\mu_{0}}.$ (5)
Since $\mu_{out}$ is the projection of $N(0,{\mathbf{\Sigma}}^{(out)}/D)$ on the unit sphere, its density is maximal at the top eigenvector of ${\mathbf{\Sigma}}^{(out)}/D$ and it is minimal at the smallest eigenvector of ${\mathbf{\Sigma}}^{(out)}/D$. Moreover,
[…]
Then the RHS of (4) is bounded as follows, while being explained below:
$\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim\mu_{out}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}\leq C_{\mu_{out},u}\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim\mu_{0}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}=C_{\mu_{out},u}\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}$ (6)
$=$ $C_{\mu_{out},u}\max\Big{(}\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{{\mathbf{U}}_{L_{*}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|},\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}^{\perp}}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}\Big{)}.$
The first inequality applies (5). The next equality replaces $\mu_{0}$ with $N(0,{\mathbf{I}})$ as the expression is scale invariant. The last equality uses the fact that $[{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}]_{L_{*},L_{*}^{\perp}}=[{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}]_{L_{*}^{\perp},L_{*}}=0$, which we clarify next. For any $\mathbf{x}\in{\mathbb{R}}^{D}$, denote $\tilde{\mathbf{x}}={\mathbf{P}}_{L_{*}}\mathbf{x}-{\mathbf{P}}_{L_{*}^{\perp}}\mathbf{x}$ and note that for $\mathbf{x}\sim N(0,{\mathbf{I}})$, $\tilde{\mathbf{x}}\sim N(0,{\mathbf{I}})$. As a result,
$2{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}={\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\Big{(}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}+\frac{\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\tilde{\mathbf{x}}\|^{2}}\Big{)}={\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{\mathbf{x}\mathbf{x}^{\top}+\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}.$
Note that $\mathbf{x}\mathbf{x}^{\top}+\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{\top}=2({\mathbf{U}}_{L_{*}}\mathbf{x})({\mathbf{U}}_{L_{*}}\mathbf{x})^{\top}+2({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})^{\top}$, so $[2({\mathbf{U}}_{L_{*}}\mathbf{x})({\mathbf{U}}_{L_{*}}\mathbf{x})^{\top}+2({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})^{\top}]_{L_{*},L_{*}}=0$ and $[2({\mathbf{U}}_{L_{*}}\mathbf{x})({\mathbf{U}}_{L_{*}}\mathbf{x})^{\top}+2({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})^{\top}]_{L_{*}^{\perp},L_{*}^{\perp}}=0$ and (6) is concluded.
Next, we bound the two terms on the RHS of (6). We claim that ${\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}^{\perp}}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}$ is a scalar matrix of size $D-d$ by $D-d$. We first note that this matrix is the same as ${\mathbb{E}}_{\mathbf{x}\sim\mu_{D-d}}\mathbf{x}\mathbf{x}^{\top}$, where $\mu_{D-d}$ denotes the uniform distribution on the unit sphere in ${\mathbb{R}}^{D-d}$. We also note that any orthogonal matrix ${\mathbf{V}}\in{\mathbb{R}}^{D-d\times D-d}$
${\mathbb{E}}_{\mathbf{x}\sim\mu_{D-d}}\mathbf{x}\mathbf{x}^{\top}={\mathbb{E}}_{\mathbf{x}\sim\mu_{D-d}}({\mathbf{V}}\mathbf{x})({\mathbf{V}}\mathbf{x})^{\top}={\mathbf{V}}({\mathbb{E}}_{\mathbf{x}\sim\mu_{D-d}}\mathbf{x}\mathbf{x}^{\top}){\mathbf{V}}^{\top}.$
Consequently, ${\mathbb{E}}_{\mathbf{x}\sim\mu_{D-d}}\mathbf{x}\mathbf{x}^{\top}$ is a scalar matrix. In other words, ${\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}^{\perp}}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}$ is a scalar matrix of size $D-d$ by $D-d$, and thus
$\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}^{\perp}}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}=\frac{1}{D-d}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\operatorname{tr}\Big{(}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}^{\perp}}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{)}=\frac{1}{D-d}.$ (7)
Since ${\mathbf{U}}_{L_{*}}^{\top}\mathbf{x}$ and ${\mathbf{U}}_{L_{*}^{\perp}}^{\top}\mathbf{x}$ are independently sampled from the Gaussian distributions $N(0,{\mathbf{I}}_{d})$ and $N(0,{\mathbf{I}}_{D-d})$,
$\Big{\|}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{{\mathbf{U}}_{L_{*}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}\Big{\|}={\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}\frac{1}{{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}}\|{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{I}})}{\mathbf{U}}_{L_{*}}^{\top}\mathbf{x}\mathbf{x}^{\top}{\mathbf{U}}_{L_{*}}\|=\frac{1}{D-d-2}\cdot 1=\frac{1}{D-d-2}.$ (8)
The last inequality uses the fact that the distribution of $\frac{1}{{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}}$ is inverse-chi-squared with parameter $D-d$ and its expectation is thus $1/(D-d-2)$.
Combining (4) and (6)-(8), we obtain the following estimate of ${\mathcal{A}}$:
$\lim_{n_{0}\rightarrow\infty}\frac{{\mathcal{A}}}{n_{0}}\leq C_{\mu_{out},u}\cdot\frac{1}{D-d-2}.$ (9)
Next, we estimate ${\mathcal{S}}$ as follows and explain the steps below:
$\lim_{n_{0}\rightarrow\infty}\frac{{\mathcal{S}}}{n_{0}}\geq\sigma_{d+1,D}\Big{(}{\mathbb{E}}_{\mathbf{x}\sim N(0,{\mathbf{\Sigma}}^{(out)}/D)}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}\Big{)}\geq\sigma_{D}\Big{(}{\mathbb{E}}_{\mathbf{x}\sim\mu_{out}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}\Big{)}\geq C_{\mu_{out},l}\sigma_{D}\Big{(}{\mathbb{E}}_{\mathbf{x}\sim\mu_{0}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}\Big{)}=\frac{C_{\mu_{out},l}}{D}.$ (10)
The first inequality follows from the definition of ${\mathcal{S}}$ and the law of large numbers, where the inequality is needed as ${\mathcal{S}}$ is defined for all data points and the RHS is for the outliers only. The second inequality just uses the definition of $\mu_{out}$ and $\sigma_{d+1,D}$. The third inequality applies (5). The last equality uses the fact that ${\mathbb{E}}_{\mathbf{x}\sim\mu_{0}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}={\mathbf{I}}/D$. To see this, we first note by symmetry that ${\mathbb{E}}_{\mathbf{x}\sim\mu_{0}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}$ is a scalar matrix and then observe that $\operatorname{tr}({\mathbb{E}}_{\mathbf{x}\sim\mu_{0}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}})=1$.
Using these estimates we can clarify Assumption 3. We recall that (3) in Assumption 3 is $\kappa_{1}\geq C\,\frac{d\,\kappa_{in}\,{\mathcal{A}}}{n_{1}}\left(\kappa_{in}+\frac{{\mathcal{A}}}{\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}}+\frac{\kappa_{2}{\mathcal{A}}}{\gamma{\mathcal{S}}}(1+\kappa_{in})\right)$, combining the estimates of ${\mathcal{A}}$ and ${\mathcal{S}}$ in (9) and (10), and the lower bound on $C$ in (62), we obtain that as $n_{0}\rightarrow\infty$, the RHS of (3) has an asymptotic upper bound of
$C\,\frac{d\,\kappa_{in}\,\kappa_{out}n_{0}}{n_{1}(D-d-2)}\left(\kappa_{in}+\kappa_{out}\frac{\frac{n_{0}}{D-d-2}}{\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}}+\frac{\kappa_{2}\kappa_{out}D}{\gamma(D-d-2)}(1+\kappa_{in})\right)$
$=$ $\max\Big{(}58,\frac{22\cdot\text{DS-SNR}-10\cdot\gamma}{\text{DS-SNR}-\gamma}\Big{)}\frac{1}{\text{DS-SNR}}\,\frac{\,\kappa_{in}\,\kappa_{out}(D-d)}{(D-d-2)}\left(\kappa_{in}+\kappa_{out}\frac{\frac{D-d}{D-d-2}}{\text{DS-SNR}-\gamma}+\frac{\kappa_{2}\kappa_{out}D}{\gamma(D-d-2)}(1+\kappa_{in})\right).$ (11)
Assuming for simplicity that $\kappa_{in},\kappa_{out}=O(1)$, which means that the inliers and outliers are sufficiently spread out in their respective spaces, and $D-d\geq 3$, then (11) can be simplified to the following order:
$\max\Big{(}58,\frac{22\cdot\text{DS-SNR}-10\cdot\gamma}{\text{DS-SNR}-\gamma}\Big{)}\,\frac{1}{\text{DS-SNR}}\left(1+\frac{1}{\text{DS-SNR}-\gamma}+\frac{1}{\gamma}\right),$ (12)
[…]
3 Proof of Main Results
We prove the main stated theorems and provide novel theoretical foundations for understanding the STE framework. In particular, our theory clarifies some statements made in Section 1.3. While the theory of TME mainly relies on its geodesic convexity, STE is not geodesically convex and we are required to develop completely novel theoretical tools, which rely on various geometric insights in the theory of positive definite matrices.
3.1 Proof of Theorem 1
The proof is organized as follows: §3.1.1 lists definitions and notation; §3.1.2 presents several technical estimates used throughout the proof.
3.1.1 Notation and Definitions
We use the common notation ${\mathbf{X}}\preceq{\mathbf{X}}^{\prime}$ for ${\mathbf{X}}$, ${\mathbf{X}}^{\prime}\in{\mathbb{R}}^{D\times D}$, which means that ${\mathbf{X}}^{\prime}-{\mathbf{X}}$ is positive semidefinite. We use $\sigma_{i}(\cdot)$ to denote the $i$-th eigenvalue of a corresponding matrix. We also use ${\mathbf{u}}_{i}(\cdot)$ to denote the $i$-th eigenvalue of a corresponding matrix. When the matrix is clear from the context, we may just write $\sigma_{i}$ and ${\mathbf{u}}_{i}$. When the matrix is positive semidefinite, then $\sigma_{i}$ is also the $i$-th singular value of this matrix, but in general this is not true.
Let $\Pi_{d}({\mathbf{\Sigma}})$ denote the projector of ${\mathbf{\Sigma}}$ onto the nearest rank-$d$ matrix, that is,
$\Pi_{d}({\mathbf{\Sigma}})=\sum_{i=1}^{d}{\mathbf{u}}_{i}({\mathbf{\Sigma}})\sigma_{i}({\mathbf{\Sigma}}){\mathbf{u}}_{i}({\mathbf{\Sigma}})^{\top}$
and ${\mathbf{P}}_{d+1,D}$ denote the projector onto the span of the smallest $D-d$ eigenvectors of ${\mathbf{\Sigma}}$, that is,
${\mathbf{P}}_{d+1,D}({\mathbf{\Sigma}})=\sum_{i=d+1}^{D}{\mathbf{u}}_{i}({\mathbf{\Sigma}}){\mathbf{u}}_{i}({\mathbf{\Sigma}})^{\top}.$
We recall our notation $\sigma_{d+1,D}({\mathbf{\Sigma}})=\frac{1}{D-d}\sum_{i=d+1}^{D}\sigma_{i}({\mathbf{\Sigma}})$ for the average of the smallest $D-d$ eigenvalues of ${\mathbf{\Sigma}}$. We further denote
$\Pi_{d+1,D}({\mathbf{\Sigma}})=\sum_{i=d+1}^{D}{\mathbf{u}}_{i}({\mathbf{\Sigma}})\sigma_{i}({\mathbf{\Sigma}}){\mathbf{u}}_{i}({\mathbf{\Sigma}})^{\top}.$
We remark that we use boldface notation for the operator ${\mathbf{P}}$, since it is linear and we identify it with its matrix representation. We do not use the boldface notation for $\Pi$ since it is not a linear operator, even though its image is a matrix.
We define the following operators from $S_{+}$ to $S_{+}$:
$T_{1}({\mathbf{\Sigma}})$ $=\sum_{{\bm{x}}\in{\mathcal{X}}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}{\mathbf{\Sigma}}^{-1}{\bm{x}}},$
$T_{2}({\mathbf{\Sigma}};\gamma)$ $=\Pi_{d}({\mathbf{\Sigma}})+\gamma\,\sigma_{d+1,D}({\mathbf{\Sigma}})\,{\mathbf{P}}_{d+1,D}({\mathbf{\Sigma}}),$
$T({\mathbf{\Sigma}})$ $=T_{2}(T_{1}({\mathbf{\Sigma}}),\gamma).$
The procedure described in Algorithm 1 is the fixed-point iteration of $T$ up to a scaling factor. We will thus explore here the equivalent subspace recovery by the fixed-point iteration of $T$.
We further define
${\mathbf{\Sigma}}_{+,in}=\sum_{{\bm{x}}\in{\mathcal{X}}_{in}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}{\mathbf{\Sigma}}^{-1}{\bm{x}}},$
${\mathbf{\Sigma}}_{+,out}=\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}{\mathbf{\Sigma}}^{-1}{\bm{x}}},$
and
${\mathbf{\Sigma}}_{+}={\mathbf{\Sigma}}_{\text{+},in}+{\mathbf{\Sigma}}_{+,out}.$
We note that
${\mathbf{\Sigma}}_{+}=T_{1}({\mathbf{\Sigma}})\ \text{ and }\ T({\mathbf{\Sigma}})=T_{2}({\mathbf{\Sigma}}_{+},\gamma).$ (13)
we define $g_{2}:S_{+}\rightarrow S_{+}$ and $g_{1}:S_{+}\rightarrow S_{+}$ by
$g_{2}({\mathbf{\Sigma}})=\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}\equiv\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}\Big{(}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}},{\mathbf{I}}\Big{)},\hskip 14.45377pt$ (14)
$g_{1}({\mathbf{\Sigma}})={\mathbf{\Sigma}}-g_{2}({\mathbf{\Sigma}})=\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}-{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&0\\ 0&0\end{pmatrix}.$ (15)
We define $g_{1}$ and $g_{2}$ for general matrices in $S_{+}$, where ${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}$ may not be invertible, by replacing ${\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}$ with $\lim_{\epsilon\rightarrow 0}{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}+\epsilon{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}$. For example,
$\text{if }\ {\mathbf{\Sigma}}=\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}&0\\ 0&0\end{pmatrix},\ \text{ then }\ g_{1}({\mathbf{\Sigma}})={\mathbf{\Sigma}}\ \text{ and }\ g_{2}({\mathbf{\Sigma}})=0.$
We recall that ${\mathbf{\Sigma}}_{in,*}$ denotes the TME solution to the set of the projected inliers $\{{\mathbf{U}}_{L^{*}}{\bm{x}}\mid{\bm{x}}\in{\mathcal{X}}_{in}\}\subset{\mathbb{R}}.$ We define
$h({\mathbf{\Sigma}})=\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}[g_{1}({\mathbf{\Sigma}})]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}),$
which corresponds to the $d$-th eigenvalue of the part of ${\mathbf{\Sigma}}$ expressed according to the basis of $L_{*}$ with appropriately scaled eigenvalues. For simplicity, we may denote
${\mathbf{\Sigma}}_{1}=g_{1}({\mathbf{\Sigma}})\ \text{ and }{\mathbf{\Sigma}}_{2}=g_{2}({\mathbf{\Sigma}}).$
For any ${\mathbf{\Sigma}}\in S_{+}$, we define
$\hat{\kappa}_{1}({\mathbf{\Sigma}})=\frac{h({\mathbf{\Sigma}}_{1})}{\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}\ \ \text{ and }\ \ \hat{\kappa}_{2}({\mathbf{\Sigma}})=\frac{\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}{\sigma_{D}({\mathbf{\Sigma}})},$
where $\hat{\kappa}_{1}$ and $\hat{\kappa}_{2}$ can obtain infinite values. We note that $\kappa_{2}=\hat{\kappa}_{2}({\mathbf{\Sigma}}^{(0)})$ and $\kappa_{1}$ is similar to $\hat{\kappa}_{1}({\mathbf{\Sigma}}^{(0)})$ as follows:
$\kappa_{1}=\frac{\sigma_{d}\Big{(}g_{1}({\mathbf{\Sigma}}^{(0)})\Big{)}}{\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}},\ \text{ whereas }\ \hat{\kappa}_{1}({\mathbf{\Sigma}}^{(0)})=\frac{\sigma_{d}\Big{(}{\mathbf{\Sigma}}_{in,*}^{-0.5}g_{1}({\mathbf{\Sigma}}^{(0)}){\mathbf{\Sigma}}_{in,*}^{-0.5}\Big{)}}{\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}}.$
For the matrix ${\mathbf{\Sigma}}_{in,*}\in{\mathbb{R}}^{d\times d}$, for example, we denote a special embedding into ${\mathbb{R}}^{D\times D}$ which agrees on its $(L_{*},L_{*})$ component (using our block decomposition) and is full rank as follows:
${\mathcal{M}}({\mathbf{\Sigma}}_{in,*})={\mathbf{U}}_{L_{*}}{\mathbf{\Sigma}}^{\prime}{\mathbf{U}}_{L_{*}}^{\top}+{\mathbf{P}}_{L_{*}^{\perp}}.$
3.1.2 Preliminaries: Some Useful Results
We formulate and prove several results that we will use in the main proof. Lemma 2 suggests several bounds on matrices that we use. Lemma 3 summarizes two basic and well-known properties of matrices. Lemma 4 is a sharper alternative to Davis-Kahan in our setting. Lemma 5 provides a bound on the lowest eigenvalue of a block matrix using information from other blocks. Lemma 6 establishes that the operator norm distances between $T({\mathbf{\Sigma}})$ (or $\Pi_{d}({\mathbf{\Sigma}}_{+})$) and ${\mathbf{\Sigma}}_{+,in}$ is bounded by a constant factor of $\|{\mathbf{\Sigma}}_{+,out}\|$. Lastly, Lemma 7 establishes the explicit formula of ${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}$ in terms of ${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}$, ${\mathbf{\Sigma}}_{L_{*},L_{*}}$ and ${\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}},$ under the condition that $\operatorname{rank}({\mathbf{\Sigma}})=d$. We write these lemmata and subsequent proofs in order.
Lemma 1.
The following properties hold:
(a) If ${\mathbf{\Sigma}}\in S_{+}$, then $g_{1}({\mathbf{\Sigma}})\in S_{+}$ and $g_{2}({\mathbf{\Sigma}})\in S_{+}$. If $\,{\mathbf{\Sigma}}\in S_{++}$, then $\sigma_{d}(g_{1}({\mathbf{\Sigma}}))>0$ and consequently ${\kappa}_{1}>0$ if ${\mathbf{\Sigma}}^{(0)}\in S_{++}$.
(b) If ${\mathbf{X}}$, ${\mathbf{\Sigma}}\in S_{+}$ and the range and domain of ${\mathbf{\Sigma}}-{\mathbf{X}}$ are contained in $L_{*}$, then $g_{1}({\mathbf{\Sigma}})\succeq{\mathbf{\Sigma}}-{\mathbf{X}}$.
Proof of Lemma 1.
(a) A direct application of [1, Theorem 1.3.3] implies that $g_{1}({\mathbf{\Sigma}})\in S_{+}$ whenever ${\mathbf{\Sigma}}\in S_{+}$. In addition, by the definition in (14),
$g_{2}({\mathbf{\Sigma}})=\begin{pmatrix}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\\ {\mathbf{I}}\end{pmatrix}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{(}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}},{\mathbf{I}}\Big{)}=\left(\begin{pmatrix}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\\ {\mathbf{I}}\end{pmatrix}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{0.5}\right)\left(\begin{pmatrix}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\\ {\mathbf{I}}\end{pmatrix}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{0.5}\right)^{T}\hskip 14.45377pt$
is also positive semidefinite.
If ${\mathbf{\Sigma}}\in S_{++}$, then (by observing the RHS of (15)),
$\operatorname{rank}(g_{2}({\mathbf{\Sigma}}))=D-d\text{ and }\ \operatorname{rank}(g_{1}({\mathbf{\Sigma}}))=d.$
Combining this with $g_{1}({\mathbf{\Sigma}})\in S_{+}$ implies that $\sigma_{d}(g_{1}({\mathbf{\Sigma}}))>0$.
If ${\mathbf{\Sigma}}^{(0)}\in S_{++}$, then ${\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}={\mathbf{U}}_{L_{*}^{\perp}}^{\top}{\mathbf{\Sigma}}^{(0)}{\mathbf{U}}_{L_{*}^{\perp}}\in S_{++}$, and thus $\kappa_{1}\equiv\sigma_{d}(g_{1}({\mathbf{\Sigma}}^{(0)}))/\sigma_{1}({\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}})>0$.
(b) Since the domain and range of ${\mathbf{\Sigma}}-{\mathbf{X}}$ is in $L_{*}$, we can express ${\mathbf{X}}$ as follows:
${\mathbf{X}}=\begin{pmatrix}{\mathbf{X}}_{L_{*},L_{*}}&{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}.$
Since ${\mathbf{X}}\in S_{+}$, [1, Theorem 1.3.3] implies that ${\mathbf{X}}_{L_{*},L_{*}}\succeq{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}$, and as a result,
$[g_{1}({\mathbf{\Sigma}})-({\mathbf{\Sigma}}-{\mathbf{X}})]_{L_{*},L_{*}}=({\mathbf{\Sigma}}_{L_{*},L_{*}}-{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}})-({\mathbf{\Sigma}}_{L_{*},L_{*}}-{\mathbf{X}}_{L_{*},L_{*}})={\mathbf{X}}_{L_{*},L_{*}}-{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\succeq 0.$
This observation, the fact that $g_{1}({\mathbf{\Sigma}})\in S_{+}$ and the above expression for ${\mathbf{X}}$ imply that $g_{1}({\mathbf{\Sigma}})\succeq{\mathbf{\Sigma}}-{\mathbf{X}}$.
Lemma 2.
The following bounds hold:
$h({\mathbf{\Sigma}}_{+,in})\geq\frac{n_{1}}{d}h({\mathbf{\Sigma}}).$ (16)
${\mathbf{\Sigma}}_{+,out}\preceq\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}$ (17)
and consequently,
$\|{\mathbf{\Sigma}}_{+,out}\|\leq{\mathcal{A}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})\ \text{ and }\ \operatorname{tr}([{\mathbf{\Sigma}}_{+,out}]_{L_{*}^{\perp},L_{*}^{\perp}})\leq{n_{0}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}).$ (18)
${\mathbf{\Sigma}}_{+,out}\succeq\sigma_{D}({\mathbf{\Sigma}})\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}},\hskip 14.45377pt{\mathbf{\Sigma}}_{+}\succeq\sigma_{D}({\mathbf{\Sigma}})\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}.$ (19)
Proof of Lemma 2.
(a) We note that $h({\mathbf{\Sigma}}_{+,in})=\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}[g_{1}({\mathbf{\Sigma}}_{+,in})]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5})=\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5})$, where we used the fact that $g_{1}({\mathbf{\Sigma}}_{+,in})={\mathbf{\Sigma}}_{+,in}$, which follows by the definition of $g_{1}$. To conclude this property, we will prove that ${\mathbf{\Sigma}}_{in,*}^{-0.5}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}\succeq h({\mathbf{\Sigma}}_{1})\frac{n_{1}}{d}{\mathbf{I}}$.
Recall that ${\mathbf{\Sigma}}_{+,in}=\sum_{{\bm{x}}\in{\mathcal{X}}_{in}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}{\mathbf{\Sigma}}^{-1}{\bm{x}}}$ and denote ${\bf{M}}={\mathcal{M}}({\mathbf{\Sigma}}_{in,*}^{-0.5})\equiv{\mathbf{U}}_{L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}{\mathbf{U}}_{L_{*}}^{\top}+{\mathbf{P}}_{L_{*}^{\perp}}$ and $\tilde{{\mathbf{X}}}_{in}=\{{\mathbf{U}}_{L_{*}}\mathbf{x}:\mathbf{x}\in{\mathcal{X}}_{in}\}\in{\mathbb{R}}^{d}$ be the representation of the set of inliers within $L_{*}$, note that
${\mathbf{\Sigma}}_{in,*}^{-0.5}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}=\sum_{{\bm{x}}\in{\mathcal{X}}_{in}}\frac{{\mathbf{\Sigma}}_{in,*}^{-0.5}[{\bm{x}}{\bm{x}}^{\top}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}}{{\bm{x}}^{\top}{\bf{M}}({\bf{M}}{\mathbf{\Sigma}}{\bf{M}})^{-1}{\bf{M}}{\bm{x}}}$
$\succeq$ $\sum_{{\bm{x}}\in{\mathcal{X}}_{in}}\frac{{\mathbf{\Sigma}}_{in,*}^{-0.5}[{\bm{x}}{\bm{x}}^{\top}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}}{{\bm{x}}^{\top}{\bf{M}}({\bf{M}}g_{1}({\mathbf{\Sigma}}){\bf{M}})^{-1}{\bf{M}}{\bm{x}}}=\sum_{\tilde{{\bm{x}}}\in\tilde{{\mathcal{X}}}_{in}}\frac{{\mathbf{\Sigma}}_{in,*}^{-0.5}\tilde{{\bm{x}}}\tilde{{\bm{x}}}^{\top}{\mathbf{\Sigma}}_{in,*}^{-0.5}}{\tilde{{\bm{x}}}^{\top}{\mathbf{\Sigma}}_{in,*}^{-0.5}({\mathbf{\Sigma}}_{in,*}^{-0.5}[g_{1}({\mathbf{\Sigma}})]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5})^{-1}\tilde{{\bm{x}}}}$
$\succeq$ $h({\mathbf{\Sigma}})\sum_{\tilde{{\bm{x}}}\in\tilde{{\mathcal{X}}}_{in}}\frac{{\mathbf{\Sigma}}_{in,*}^{-0.5}\tilde{{\bm{x}}}\tilde{{\bm{x}}}^{\top}{\mathbf{\Sigma}}_{in,*}^{-0.5}}{\tilde{{\bm{x}}}^{\top}{\mathbf{\Sigma}}_{in,*}^{-1}\tilde{{\bm{x}}}}h({\mathbf{\Sigma}})\frac{n_{1}}{d}{\mathbf{I}}.$
The first matrix inequality follows from the fact that $g_{1}({\mathbf{\Sigma}})\preceq{\mathbf{\Sigma}}$ (which is obvious by the definition of $g_{1}$). Indeed, this inequality implies that ${\mathbf{\Sigma}}_{in,*}^{-0.5}g_{1}({\mathbf{\Sigma}}){\mathbf{\Sigma}}_{in,*}^{-0.5}\preceq{\mathbf{\Sigma}}_{in,*}^{-0.5}{\mathbf{\Sigma}}{\mathbf{\Sigma}}_{in,*}^{-0.5}$, which leads to $({\mathbf{\Sigma}}_{in,*}^{-0.5}g_{1}({\mathbf{\Sigma}}){\mathbf{\Sigma}}_{in,*}^{-0.5})^{-1}\succeq({\mathbf{\Sigma}}_{in,*}^{-0.5}{\mathbf{\Sigma}}{\mathbf{\Sigma}}_{in,*}^{-0.5})^{-1}$ and consequently $\mathbf{x}^{\top}({\mathbf{\Sigma}}_{in,*}^{-0.5}g_{1}({\mathbf{\Sigma}}){\mathbf{\Sigma}}_{in,*}^{-0.5})^{-1}\mathbf{x}\geq\mathbf{x}^{\top}({\mathbf{\Sigma}}_{in,*}^{-0.5}{\mathbf{\Sigma}}{\mathbf{\Sigma}}_{in,*}^{-0.5})^{-1}\mathbf{x}$.
The second equality follows from the fact that $g_{1}({\mathbf{\Sigma}})$ is a singular matrix with range $L_{*}$, and $\mathbf{x}\in{\mathcal{X}}_{in}$ lie in $L_{*}$.
The second inequality similarly follows from the fact ${\mathbf{\Sigma}}_{in,*}^{-0.5}[g_{1}({\mathbf{\Sigma}})]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}\succeq h({\mathbf{\Sigma}}){\mathbf{I}}$, which is a direct consequence of the definition $h({\mathbf{\Sigma}}):=\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}[g_{1}({\mathbf{\Sigma}})]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5})$.
The last equality applies the assumption that ${\mathbf{\Sigma}}_{in,*}$ is well-defined as a TME solution to $\tilde{{\mathcal{X}}}$ (which follows from Assumption 1 and Section 4 of [4]), and the following property of TME: if ${\mathbf{\Sigma}}^{\text{TME}}$ is a TME solution to a set ${\mathcal{X}}\subset{\mathbb{R}}^{p}$, then
$\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}^{\top}\mathbf{x}}{\mathbf{x}^{\top}({\mathbf{\Sigma}}^{\text{TME}})^{-1}\mathbf{x}}=\frac{|{\mathcal{X}}|}{p}{\mathbf{\Sigma}}^{\text{TME}},$ (20)
where for our case $p=d$ and $|\tilde{{\mathcal{X}}_{in}}|=n_{1}$.
(b) Equation (17) is proved as follows (where we explain the details of proof below):
${\mathbf{\Sigma}}_{+,out}$ $=\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}{\mathbf{\Sigma}}^{-1}{\bm{x}}}\preceq\lim_{t\rightarrow\infty}\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}(t{\mathbf{P}}_{L_{*}}+{\mathbf{\Sigma}})^{-1}{\bm{x}}}=\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{\mathbf{x}^{T}{\mathbf{U}}_{L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-1}{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}}$ (21)
$\preceq$ $\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{\mathbf{x}^{T}{\mathbf{U}}_{L_{*}^{\perp}}\frac{{\mathbf{I}}}{\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}}=\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{\|{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\|^{2}}.$
The first equality follows from the definition of ${\mathbf{\Sigma}}_{+,out}$. The first inequality follows from the fact that $t{\mathbf{P}}_{L_{*}}+{\mathbf{\Sigma}}\succeq{\mathbf{\Sigma}}$ for $t>0$ and thus $\mathbf{x}^{T}{\mathbf{\Sigma}}^{-1}\mathbf{x}\geq\mathbf{x}^{T}(t{\mathbf{P}}_{L_{*}}+{\mathbf{\Sigma}})^{-1}\mathbf{x}$. The second equality follows from observing that $\lim_{t\rightarrow\infty}(t{\mathbf{P}}_{L_{*}}+{\mathbf{\Sigma}})^{-1}={\mathbf{U}}_{L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-1}{\mathbf{U}}_{L_{*}^{\perp}}^{T}$. Indeed, by letting $t\to\infty$, the 3 other blocks, associated with $(L_{*},L_{*})$, $(L_{*},L_{*}^{\perp})$ and $(L_{*}^{\perp},L_{*})$ zero out in the inverse and the remaining one, that is the one associated with $(L_{*}^{\perp},L_{*}^{\perp})$, remains the same in the inverse.
Lastly, (18) is concluded by applying the norm and trace to both sides of (17).
(c) The proof of (19) is straightforward and uses the fact that ${\mathbf{\Sigma}}\succeq\sigma_{D}({\mathbf{\Sigma}}){\mathbf{I}}$. We exemplify it for bounding from below ${\mathbf{\Sigma}}_{+,out}$, where the proof for ${\mathbf{\Sigma}}_{+}$ is the same:
${\mathbf{\Sigma}}_{+,out}=\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}{\mathbf{\Sigma}}^{-1}{\bm{x}}}\succeq\sum_{{\bm{x}}\in{\mathcal{X}}_{out}}\frac{{\bm{x}}{\bm{x}}^{\top}}{{\bm{x}}^{\top}(\sigma_{D}({\mathbf{\Sigma}}){\mathbf{I}})^{-1}{\bm{x}}}=\sigma_{D}({\mathbf{\Sigma}})\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}.$
Lemma 3.
The following two properties hold:
(a) For any semi-orthogonal matrix ${\mathbf{U}}\in O(D,d)$ and symmetric matrix ${\mathbf{\Sigma}}\in{\mathbb{R}}^{D\times D}$,
[…]
(b) For any two symmetric matrices ${\mathbf{A}},{\mathbf{B}}$, $|\sigma_{i}({\mathbf{A}})-\sigma_{i}({\mathbf{B}})|\leq\|{\mathbf{A}}-{\mathbf{B}}\|$.
The proof of part (a) follows from the Courant-Fischer min-max theorem for the eigenvalues [9, Theorem 1.3.2]. For completeness we verify this as follows:
$\sigma_{D-d+1}({\mathbf{\Sigma}})=\min_{\dim(V)=d}\ \max_{{\mathbf{v}}\in V:\|{\mathbf{v}}\|=1}{\mathbf{v}}^{T}{\mathbf{\Sigma}}{\mathbf{v}}=\min_{{\mathbf{V}}\in O(D,d)}\ \max_{{\mathbf{u}}\in{\mathbb{R}}^{d}:\|{\mathbf{u}}\|=1}{\mathbf{u}}^{T}{\mathbf{V}}^{T}{\mathbf{\Sigma}}{\mathbf{V}}{\mathbf{u}}=\min_{{\mathbf{V}}\in O(D,d)}\sigma_{1}({\mathbf{V}}^{T}{\mathbf{\Sigma}}{\mathbf{V}})\leq\sigma_{1}({\mathbf{U}}^{T}{\mathbf{\Sigma}}{\mathbf{U}})$
and the part $\sigma_{1}({\mathbf{\Sigma}})\geq\sigma_{1}({\mathbf{U}}^{T}{\mathbf{\Sigma}}{\mathbf{U}})$ follows from $\sigma_{1}({\mathbf{\Sigma}})=\max_{{\mathbf{v}}:\|{\mathbf{v}}\|=1}{\mathbf{v}}^{T}{\mathbf{\Sigma}}{\mathbf{v}}$.
On the other hand,
$\sigma_{d}({\mathbf{\Sigma}})=\max_{\dim(V)=d}\ \min_{{\mathbf{v}}\in V:\|{\mathbf{v}}\|=1}{\mathbf{v}}^{T}{\mathbf{\Sigma}}{\mathbf{v}}=\max_{{\mathbf{V}}\in O(D,d)}\ \min_{{\mathbf{u}}\in{\mathbb{R}}^{d}:\|{\mathbf{u}}\|=1}{\mathbf{u}}^{T}{\mathbf{V}}^{T}{\mathbf{\Sigma}}{\mathbf{V}}{\mathbf{u}}=\max_{{\mathbf{V}}\in O(D,d)}\sigma_{d}({\mathbf{V}}^{T}{\mathbf{\Sigma}}{\mathbf{V}})\geq\sigma_{d}({\mathbf{U}}^{T}{\mathbf{\Sigma}}{\mathbf{U}})$
and the part $\sigma_{d}({\mathbf{U}}^{T}{\mathbf{\Sigma}}{\mathbf{U}})\geq\sigma_{D}({\mathbf{\Sigma}})$ follows from $\sigma_{D}({\mathbf{\Sigma}})=\min_{{\mathbf{v}}:\|{\mathbf{v}}\|=1}{\mathbf{v}}^{T}{\mathbf{\Sigma}}{\mathbf{v}}$.
For part (b) see (1.63) of [9] and its proof in there.
Lemma 4.
If ${\mathbf{\Sigma}}\in S_{+}(D)$ and $\hat{L}$ is the span of the top $d$ eigenvectors of ${\mathbf{\Sigma}}$, then
$\sin\angle(\hat{L},L_{*})\leq 2\sqrt{\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})/\sigma_{d}({\mathbf{\Sigma}}_{L_{*},L_{*}})}.$
Proof of Lemma 4.
For ${\mathbf{\Sigma}}\in S_{+}$, we denote $\kappa_{0}=\sigma_{d}({\mathbf{\Sigma}}_{L_{*},L_{*}})/\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})$. Using this notation, we need to prove that
$\sin\angle(\hat{L},L_{*})\leq 2/\sqrt{\kappa_{0}}.$
Assume on the contrary that this is not true and thus $\sin\angle(\hat{L}^{\perp},L_{*}^{\perp})=\sin\angle(\hat{L},L_{*})>2/\sqrt{\kappa_{0}}$. That is, there exists a unit vector ${\mathbf{v}}\in\hat{L}^{\perp}$ such that $\sin\angle({\mathbf{v}},L_{*}^{\perp})>2/\sqrt{\kappa_{0}}$. We express this vector ${\mathbf{v}}\in\hat{L}^{\perp}$ as follows:
${\mathbf{v}}=\sin\theta{\mathbf{v}}_{1}+\cos\theta{\mathbf{v}}_{2},$
where ${\mathbf{v}}_{1}\in L_{*}$ and ${\mathbf{v}}_{2}\in L_{*}^{\perp}$. Our assumption then implies that $\sin\theta>2/\sqrt{\kappa_{0}}$. For simplicity, we assume without loss of generality that $\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})=1$ and thus $\sigma_{d}({\mathbf{\Sigma}}_{L_{*},L_{*}})=\kappa_{0}$. We thus obtain that ${\mathbf{v}}_{1}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{1}\geq\kappa_{0}$ and ${\mathbf{v}}_{2}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{2}\leq 1$. Since ${\mathbf{\Sigma}}$ is p.s.d., $|{\mathbf{v}}_{1}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{2}|\leq\sqrt{{\mathbf{v}}_{1}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{1}}\sqrt{{\mathbf{v}}_{2}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{2}}$. Consequently,
${\mathbf{v}}^{T}{\mathbf{\Sigma}}{\mathbf{v}}$
$=\sin^{2}\theta{\mathbf{v}}_{1}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{1}+\cos^{2}\theta{\mathbf{v}}_{2}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{2}-2\sin\theta\cos\theta{\mathbf{v}}_{1}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{2}$
$\geq\left(\sin\theta\sqrt{{\mathbf{v}}_{1}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{1}}-\cos\theta\sqrt{{\mathbf{v}}_{2}^{T}{\mathbf{\Sigma}}{\mathbf{v}}_{2}}\right)^{2}$
$\geq(\sin\theta\sqrt{\kappa_{0}}-\cos\theta)^{2}$
$>(2-1)^{2}=1.$
We show on the other hand that ${\mathbf{v}}^{T}{\mathbf{\Sigma}}{\mathbf{v}}\leq 1$ and thus obtain a contradiction, which concludes the proof. Indeed, since ${\mathbf{v}}\in\hat{L}^{\perp}$, it is spanned by the smallest $D-d$ eigenvectors of ${\mathbf{\Sigma}}$, and consequently ${\mathbf{v}}^{T}{\mathbf{\Sigma}}{\mathbf{v}}\leq\sigma_{d+1}({\mathbf{\Sigma}})\leq\sigma_{1}({\mathbf{U}}_{L_{*}^{\perp}}^{\top}{\mathbf{\Sigma}}{\mathbf{U}}_{L_{*}^{\perp}})=1$, where the second inequality follows from Lemma 3(a)).
Lemma 5.
Assume that ${\mathbf{\Sigma}}$ is positive semidefinite and denote $x=\sigma_{D-d}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})$, $y=\sigma_{d}({\mathbf{\Sigma}}_{L_{*},L_{*}})$ and $z=\|{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}\|$. Then,
$\sigma_{D}({\mathbf{\Sigma}})\geq\frac{(x+y)-\sqrt{(x-y)^{2}+4z^{2}}}{2}$ (22)
and
$\text{if }\ z\leq\sqrt{xy}/2,\ \text{ then }\ \sigma_{D}({\mathbf{\Sigma}})\geq\min(x,y)/3.$ (23)
Proof of Lemma 5.
Let $a=\frac{(x+y)-\sqrt{(x-y)^{2}+4z^{2}}}{2}$ and note that
$a=\frac{(x+y)-\sqrt{(x-y)^{2}+4z^{2}}}{2}\leq\frac{(x+y)-\sqrt{(x-y)^{2}}}{2}\leq\frac{(x+y)-(y-x)}{2}=x.$
It thus follows from the definition of $x$ and this inequality that
$({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})\in S_{+}.$
We also note that
$x-\frac{z^{2}}{y-a}=x-\frac{2z^{2}}{(y-x)+\sqrt{(x-y)^{2}+4z^{2}}}=x-\frac{-(y-x)+\sqrt{(x-y)^{2}+4z^{2}}}{2}=a.$ (24)
Let
${\mathbf{\Sigma}}_{0}=\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}}&{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{T},\end{pmatrix}$
so that
${\mathbf{\Sigma}}-{\mathbf{\Sigma}}_{0}=\begin{pmatrix}a{\mathbf{I}}&0\\ 0&{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}-{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{T}.\end{pmatrix}.$
We note that the eigenvalues of ${\mathbf{\Sigma}}-{\mathbf{\Sigma}}_{0}$ consist of $a$ (with multiplicity $d$) and the eigenvalues of
${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}-{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{T},$
which are bounded below as follows
$\sigma_{D-d}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}-{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{T})\geq\sigma_{D-d}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})-\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{T}))$
$\geq$ $x-\frac{\|{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\|^{2}}{\sigma_{d}({\mathbf{\Sigma}}_{L_{*},L_{*}})-a}\geq x-\frac{z^{2}}{y-a}=a.$
The above first inequality uses part (b) of Lemma 3 and the facts that ${\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}}\in S_{+}$, thus ${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{T}\in S_{+}$, and consequently the largest eigenvalue of the latter matrix coincides with its norm. The second inequality follows from basic observations, in particular, ${\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}}\preceq{\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{T})}\;{\mathbf{I}}$. The last equality is (24).
We have thus showed that all eigenvalue of ${\mathbf{\Sigma}}-{\mathbf{\Sigma}}_{0}$ are bounded below by $a$. We also note that ${\mathbf{\Sigma}}_{0}\in S_{+}$ as follows:
${\mathbf{\Sigma}}_{0}=\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\end{pmatrix}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-1}\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}}&{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}.\end{pmatrix}$
$=$ $\Bigg{(}\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\end{pmatrix}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-0.5}\Bigg{)}\Bigg{(}\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}\end{pmatrix}({\mathbf{\Sigma}}_{L_{*},L_{*}}-a{\mathbf{I}})^{-0.5}\Bigg{)}^{T}.$
Consequently, all eigenvalues of ${\mathbf{\Sigma}}$ are bounded below by $a$, that is, $\sigma_{D}({\mathbf{\Sigma}})\geq a$ and thus we proved (22).
If $z\leq\sqrt{xy}/2\leq\max(x,y)/2$, then
$a=\frac{2xy-2z^{2}}{(x+y)+\sqrt{(x-y)^{2}+4z^{2}}}\geq\frac{2xy-2(\sqrt{xy}/2)^{2}}{(x+y)+|x-y|+2z}\geq\frac{xy}{2\max(x,y)+2z}\geq\frac{xy}{2\max(x,y)+\max(x,y)}=\frac{\min(x,y)}{3},$
where the first inequality follows from $\sqrt{a_{1}^{2}+a_{2}^{2}}\leq a_{1}+a_{2}$.
Lemma 6.
For any ${\mathbf{\Sigma}}\in S_{+}$,
$\|T({\mathbf{\Sigma}})-{\mathbf{\Sigma}}_{+,in}\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|.$ (25)
and
$\|\Pi_{d}({\mathbf{\Sigma}}_{+})-{\mathbf{\Sigma}}_{+,in}\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|.$ (26)
Proof of Lemma 6.
We first note that
$\|T({\mathbf{\Sigma}})-{\mathbf{\Sigma}}_{+}\|=\|T_{2}({\mathbf{\Sigma}}_{+},\gamma)-{\mathbf{\Sigma}}_{+}\|=\|\Pi_{d}({\mathbf{\Sigma}}_{+})+\gamma\,\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\,{\mathbf{P}}_{d+1,D}({\mathbf{\Sigma}}_{+})-{\mathbf{\Sigma}}_{+}\|\\ =\|\gamma\,\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\,{\mathbf{P}}_{d+1,D}({\mathbf{\Sigma}}_{+})-\Pi_{d+1,D}({\mathbf{\Sigma}}_{+})\|\leq\sigma_{d+1}({\mathbf{\Sigma}}_{+}),$
where the last inequality follows from the fact that the nonzero eigenvalues of $\gamma\,\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\,{\mathbf{P}}_{d+1,D}({\mathbf{\Sigma}}_{+})-\Pi_{d+1,D}({\mathbf{\Sigma}}_{+})$ are $\{\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})-\sigma_{i}({\mathbf{\Sigma}}_{+})\}_{d+1\leq i\leq D}$ and they satisfy:
$\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})-\sigma_{i}({\mathbf{\Sigma}}_{+})\leq\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\leq\sigma_{d+1}({\mathbf{\Sigma}}_{+})\ \text{ and }\ \gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})-\sigma_{i}({\mathbf{\Sigma}}_{+})\geq-\sigma_{i}({\mathbf{\Sigma}}_{+})\geq-\sigma_{d+1}({\mathbf{\Sigma}}_{+}).$
We also apply Lemma 3(a) and note that
$\sigma_{d+1}({\mathbf{\Sigma}}_{+})\leq\sigma_{1}([{\mathbf{\Sigma}}_{+}]_{L_{*}^{\perp},L_{*}^{\perp}})=\sigma_{1}([{\mathbf{\Sigma}}_{+,out}]_{L_{*}^{\perp},L_{*}^{\perp}}))\leq\sigma_{1}({\mathbf{\Sigma}}_{+,out}).$
Applying twice the triangle inequality and the above two equations, we obtain the desired estimate as follows:
$\|T({\mathbf{\Sigma}})-{\mathbf{\Sigma}}_{+,in}\|\leq\|T({\mathbf{\Sigma}})-{\mathbf{\Sigma}}_{+}\|+\|{\mathbf{\Sigma}}_{+}-{\mathbf{\Sigma}}_{+,in}\|\leq\sigma_{d+1}({\mathbf{\Sigma}}_{+})+\|{\mathbf{\Sigma}}_{+,out}\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|,$
Equation 26 is similarly obtained:
$\|\Pi_{d}({\mathbf{\Sigma}}_{+})-{\mathbf{\Sigma}}_{+,in}\|\leq\sigma_{d+1}({\mathbf{\Sigma}}_{+})+\|{\mathbf{\Sigma}}_{+}-{\mathbf{\Sigma}}_{+,in}\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|.$ (27)
Lemma 7.
If ${\mathbf{\Sigma}}\in S_{+}$, $\operatorname{rank}({\mathbf{\Sigma}})=d$ and ${\mathbf{\Sigma}}_{L_{*},L_{*}}$ is invertible, then
${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}={\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}{\mathbf{\Sigma}}_{L_{*},L_{*}}^{-1}{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}.$
Proof of Lemma 7.
Since ${\mathbf{\Sigma}}\in S_{+}$ and $\operatorname{rank}({\mathbf{\Sigma}})=d$, ${\mathbf{\Sigma}}={\mathbf{U}}{\mathbf{U}}^{\top}$ with ${\mathbf{U}}\in{\mathbb{R}}^{D\times d}$. Without loss of generality, we assume that $L_{*}$ spans the first $d$ column vectors of ${\mathbf{U}}$. We write ${\mathbf{U}}=[{\mathbf{U}}_{1};{\mathbf{U}}_{2}]$ with ${\mathbf{U}}_{1}\in{\mathbb{R}}^{d\times d}$ and ${\mathbf{U}}_{2}\in{\mathbb{R}}^{(D-d)\times d}$, so this assumption means that ${\mathbf{U}}_{1}{\mathbf{U}}_{1}^{\top}={\mathbf{\Sigma}}_{L_{*},L_{*}}$ and consequently ${\mathbf{U}}_{1}={\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{0.5}$. Since ${\mathbf{U}}_{2}{\mathbf{U}}_{1}^{\top}={\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}$, we deduce that ${\mathbf{U}}_{2}={\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}^{-0.5}$. Noting further that ${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}={\mathbf{U}}_{2}{\mathbf{U}}_{2}^{\top}$, we conclude the lemma. ∎
Lemma 8.
If ${\mathbf{A}}$, ${\mathbf{B}}\in S_{+}(d)$, then
$\sigma_{d}({\mathbf{A}}{\mathbf{B}}{\mathbf{A}})\geq\sigma_{d}^{2}({\mathbf{A}})\sigma_{d}({\mathbf{B}})\ \text{ and }\ \sigma_{d}({\mathbf{A}}{\mathbf{B}}{\mathbf{A}})\leq\sigma_{1}^{2}({\mathbf{A}})\sigma_{d}({\mathbf{B}}).$
Proof of Lemma 8.
The first inequality follows from the fact that
for $\text{for }\ {\mathbf{A}}\in S_{+}(d),\ \sigma_{d}({\mathbf{A}})=\min_{\mathbf{x}\in{\mathbb{R}}^{d}}\|{\mathbf{A}}\mathbf{x}\|/\|\mathbf{x}\|.$ (28)
If ${\mathbf{A}}$ is singular, $\sigma_{d}({\mathbf{A}}{\mathbf{B}}{\mathbf{A}})=0$ and the inequality is trivial. Otherwise, we form the vector $\bar{{\mathbf{u}}}:={\mathbf{A}}^{-1}{\mathbf{u}}_{d}({\mathbf{B}})$ and note that
$\bar{{\mathbf{u}}}^{\top}({\mathbf{A}}{\mathbf{B}}{\mathbf{A}})\bar{{\mathbf{u}}}=\sigma_{d}({\mathbf{B}})\ \text{ and }\ \|\bar{{\mathbf{u}}}\|\geq\frac{1}{\sigma_{1}({\mathbf{A}})},$
where the inequality above follows from (28) applied to ${\mathbf{A}}^{-1}$, where $\sigma_{d}({\mathbf{A}}^{-1})=\sigma_{1}({\mathbf{A}})^{-1}$. Consequently,
$\sigma_{d}\big{(}{\mathbf{A}}{\mathbf{B}}{\mathbf{A}}\big{)}\leq\frac{\bar{{\mathbf{u}}}^{\top}{\mathbf{A}}{\mathbf{B}}{\mathbf{A}}\bar{{\mathbf{u}}}}{\|\bar{{\mathbf{u}}}\|^{2}}\leq\sigma_{1}^{2}({\mathbf{A}})\sigma_{d}({\mathbf{B}}).$
3.1.3 Reduction of the theorem
We reduce the proof of the theorem to proving another statement. We recall that Assumption 3 requires a sufficiently large constant $C$, so (3) is satisfied. Using this constant we define
$\tilde{\kappa}_{1}=C\,\frac{d\,{\mathcal{A}}}{n_{1}\,\sigma_{d}({\mathbf{\Sigma}}_{in,*})}\left(\kappa_{in}+\frac{{\mathcal{A}}}{\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}}+\frac{\kappa_{2}{\mathcal{A}}}{\gamma{\mathcal{S}}}(1+\kappa_{in})\right),$
We note that (3) of Assumption 3 requires that
$\kappa_{1}>\sigma_{1}({\mathbf{\Sigma}}_{in,*})\cdot\tilde{\kappa}_{1}.$
We also define
$C_{0}=2\frac{\frac{n_{1}}{d}}{\frac{n_{1}}{d}+\gamma\frac{n_{0}}{D-d}}$ (29)
and note that Assumption 2 implies that $C_{0}>1$. Using these constants and notation and definitions from §3.1.1, we formulate the alternative statement as follows:
$\text{If }\hat{\kappa}_{1}({\mathbf{\Sigma}})\geq\tilde{\kappa}_{1}\text{ and }\hat{\kappa}_{2}({\mathbf{\Sigma}})\leq\max(\kappa_{2},7),\text{ then }\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq C_{0}\hat{\kappa}_{1}({\mathbf{\Sigma}})\text{ and }\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))\leq\max(\kappa_{2},7).$ (30)
We will show next that (30) implies the conclusion of the theorem.
We first note that $\hat{\kappa}_{1}({\mathbf{\Sigma}}^{(0)})>\tilde{\kappa}_{1}$. Indeed, direct application of Lemma 8 yields
$\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}{\mathbf{\Sigma}}^{(0)}{\mathbf{\Sigma}}_{in,*}^{-0.5})\geq\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5})\sigma_{d}({\mathbf{\Sigma}}^{(0)})\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5})=\frac{\sigma_{d}({\mathbf{\Sigma}}^{(0)})}{\sigma_{1}({\mathbf{\Sigma}}_{in,*})}$
and consequently
$\hat{\kappa}_{1}({\mathbf{\Sigma}}^{(0)})=\frac{\sigma_{d}\Big{(}{\mathbf{\Sigma}}_{in,*}^{-0.5}g_{1}({\mathbf{\Sigma}}^{(0)}){\mathbf{\Sigma}}_{in,*}^{-0.5}\Big{)}}{\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}}\geq\frac{\sigma_{d}\Big{(}g_{1}({\mathbf{\Sigma}}^{(0)})\Big{)}}{\sigma_{1}({\mathbf{\Sigma}}_{in,*})\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}}=\frac{\kappa_{1}}{\sigma_{1}({\mathbf{\Sigma}}_{in,*})}>\tilde{\kappa}_{1}.$ (31)
Furthermore, since $\hat{\kappa}_{2}({\mathbf{\Sigma}}^{(0)})=\kappa_{2}$, both conditions of (30) are satisfied and we can assume the conclusion of (30) for ${\mathbf{\Sigma}}^{(k+1)}=T({\mathbf{\Sigma}}^{(k)})$, $k\geq 0$.
We note that (30) and (31) imply that $\hat{\kappa}_{1}({\mathbf{\Sigma}}^{(k)})\rightarrow\infty$ and the convergence to infinity is linear. We further note that
${\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}=({\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-0.5}_{L_{*}^{\perp},L_{*}^{\perp}})({\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-0.5}_{L_{*}^{\perp},L_{*}^{\perp}})^{T}\,\in S_{+}$
and thus
${\mathbf{\Sigma}}_{L_{*},L_{*}}\succeq g_{1}({\mathbf{\Sigma}})\equiv\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}-{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}{\mathbf{\Sigma}}^{-1}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&0\\ 0&0\end{pmatrix}$
and consequently,
$\sigma_{d}({\mathbf{\Sigma}}_{L_{*},L_{*}})/\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})\geq\sigma_{d}(g_{1}({\mathbf{\Sigma}}))/\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})=\hat{\kappa}_{1}({\mathbf{\Sigma}}).$ (32)
Combining (32) with Lemma 4, we conclude that $\sin\angle({L}^{(k)},L_{*})\leq 2/\sqrt{\hat{\kappa}_{1}({\mathbf{\Sigma}}^{(k)})}$. This observation and the fact $\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq C_{0}\hat{\kappa}_{1}({\mathbf{\Sigma}})$ (according to (30)) imply that $\angle({L}^{(k)},L_{*})$ converges $r$-linearly to zero.
Next, we show that ${\mathbf{\Sigma}}^{(k)}$ converges to a singular matrix. We first observe that
$\sigma_{1}([{\mathbf{\Sigma}}^{(k)}]_{L_{*}^{\perp},L_{*}^{\perp}})=\frac{\sigma_{1}([{\mathbf{\Sigma}}^{(k)}]_{L_{*}^{\perp},L_{*}^{\perp}})}{\operatorname{tr}({\mathbf{\Sigma}}^{(k)})}=\frac{\sigma_{1}([{\mathbf{\Sigma}}^{(k)}]_{L_{*}^{\perp},L_{*}^{\perp}})}{\operatorname{tr}([{\mathbf{\Sigma}}^{(k)}]_{L_{*},L_{*}})+\operatorname{tr}([{\mathbf{\Sigma}}^{(k)}]_{L_{*}^{\perp},L_{*}^{\perp}})}\leq\frac{\sigma_{1}([{\mathbf{\Sigma}}^{(k)}]_{L_{*}^{\perp},L_{*}^{\perp}})}{\sigma_{d}({\mathbf{\Sigma}}^{(k)}_{L_{*},L_{*}})+\sigma_{1}({\mathbf{\Sigma}}^{(k)}_{L_{*}^{\perp},L_{*}^{\perp}})}\leq\frac{1}{\hat{\kappa}_{1}({\mathbf{\Sigma}}^{(k)})+1},$
where the first equality uses the fact that STE algorithm scales ${\mathbf{\Sigma}}^{(k)}$ by its trace so $\operatorname{tr}({\mathbf{\Sigma}}^{(k)})=1$ and the last inequality follows from (32). Applying first Lemma 3(a) and then the above estimate, we conclude that
$\sigma_{d+1}({\mathbf{\Sigma}}^{(k)})\leq\sigma_{1}([{\mathbf{\Sigma}}^{(k)}]_{L_{*}^{\perp},L_{*}^{\perp}})\leq\frac{1}{\hat{\kappa}_{1}({\mathbf{\Sigma}}^{(k)})+1}\rightarrow 0\hskip 7.22743pt\text{as}\hskip 7.22743ptk\rightarrow\infty.$
That is, ${\mathbf{\Sigma}}^{(k)}$ converges to a singular matrix.
Since ${\mathbf{\Sigma}}^{(k)}$ converges to a singular matrix, $\sigma_{d+1,D}(T_{1}({\mathbf{\Sigma}}^{(k)}))$ converges to zero. Applying this observation and then the fact that $\angle({L}^{(k)},L_{*})\rightarrow 0$, we obtain that for sufficiently large $k$
${\mathbf{\Sigma}}^{(k+1)}=T_{2}(T_{1}({\mathbf{\Sigma}}^{(k)});\gamma)\equiv\Pi_{d}(T_{1}({\mathbf{\Sigma}}^{(k)}))+\gamma\,\sigma_{d+1,D}(T_{1}({\mathbf{\Sigma}}^{(k)}))\,{\mathbf{P}}_{d+1,D}(T_{1}({\mathbf{\Sigma}}^{(k)}))\to\approx\Pi_{d}(T_{1}({\mathbf{\Sigma}}^{(k)}))\approx{\mathbf{P}}_{L_{*}}T_{1}({\mathbf{\Sigma}}^{(k)}){\mathbf{P}}_{L_{*}}.$
Recall that $T_{1}$ is the standard TME procedure and thus note that the above argument implies that the STE algorithm converges to the TME algorithm restricted to be from $L_{*}$ to $L_{*}$. Since $T_{1}({\mathbf{\Sigma}}^{(k)})$ is a weighted sum of $\mathbf{x}\mathbf{x}^{T}$ with weight $1/(\mathbf{x}^{T}{\mathbf{\Sigma}}^{(k)\,-1}\mathbf{x})$, and for $\mathbf{x}\in{\mathcal{X}}_{out}$ this weight goes to 00 (since the domain and range of ${\mathbf{\Sigma}}^{(k)}$ converge to $L_{*}$),
${\mathbf{\Sigma}}^{(k+1)}\approx{{\mathbf{P}}_{L_{*}}T_{1}({\mathbf{\Sigma}}^{(k)}){\mathbf{P}}_{L_{*}}}\approx{\mathbf{P}}_{L_{*}}\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}{\mathbf{\Sigma}}^{(k)\,-1}\mathbf{x}}{\mathbf{P}}_{L_{*}}.$
We also note that the iterative update formula of TME restricted to the projection of ${\mathcal{X}}_{in}$ in $L_{*}$, that is, to $\tilde{{\mathcal{X}}}_{in}=\{{\mathbf{U}}_{L_{*}}\mathbf{x}:\mathbf{x}\in{\mathcal{X}}_{in}\}\in{\mathbb{R}}^{d}$ is
${\mathbf{\Sigma}}^{(k+1)}_{\tilde{{\mathcal{X}}}_{in}}=\sum_{\tilde{\mathbf{x}}\in\tilde{{\mathcal{X}}}_{in}}\frac{\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{T}}{\tilde{\mathbf{x}}^{T}{\mathbf{\Sigma}}^{(k)\,-1}_{\tilde{{\mathcal{X}}}_{in}}\tilde{\mathbf{x}}}.$
We thus proved that $[{\mathbf{\Sigma}}^{(k)}_{L_{*},L_{*}}]$ converges the TME solution of the inliers ${\mathbf{\Sigma}}_{*,in}$. We also proved above the $r$-linear convergence to 0 of $\angle({L}^{(k)},L_{*})$, so we verified Theorem 1 assuming (30) holds. It remains to prove (30).
3.1.4 Proof of (30)
We separately prove the two implications of (30).
Proof that $\bm{\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq C_{0}\hat{\kappa}_{1}({\mathbf{\Sigma}})}$: Recall that
$\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))=\frac{h({T}({\mathbf{\Sigma}}))}{\sigma_{1}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})},\ \text{ where }\ h(T({\mathbf{\Sigma}}))=\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}[g_{1}(T({\mathbf{\Sigma}}))]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}).$ (33)
We estimate the denominator of $\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))$ under a special condition, then its numerator, and at last use these estimates to conclude that $\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq C_{0}\hat{\kappa}_{1}({\mathbf{\Sigma}})$.
Estimate of the denominator of $\bm{\hat{\kappa}_{1}}$ under a special condition: We first obtain lower and upper bounds for $h({\mathbf{\Sigma}}_{+,in})$. Since $g_{1}({\mathbf{\Sigma}}_{+,in})={\mathbf{\Sigma}}_{+,in}$,
$h({\mathbf{\Sigma}}_{+,in})=\sigma_{d}\big{(}{\mathbf{\Sigma}}_{in,*}^{-0.5}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}\big{)}\geq\sigma_{d}^{2}\big{(}{\mathbf{\Sigma}}_{in,*}^{-0.5}\big{)}\sigma_{d}\big{(}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}})\sigma_{d}\big{(}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}\big{)}/\sigma_{1}({\mathbf{\Sigma}}_{in,*}),$ (34)
where the inequality in the above equation follows from Lemma 8 and the fact that the matrices ${\mathbf{\Sigma}}_{in,*}^{-0.5}$ and $[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}$ are in $S_{+}(d)$.
To bound $h({\mathbf{\Sigma}}_{+,in})$ from above, we again apply Lemma 8 and obtain the following bounds:
$h({\mathbf{\Sigma}}_{+,in})=\sigma_{d}\big{(}{\mathbf{\Sigma}}_{in,*}^{-0.5}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5}\big{)}\leq\frac{\sigma_{d}\big{(}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}\big{)}}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}=\frac{\sigma_{d}\big{(}{\mathbf{\Sigma}}_{+,in}\big{)}}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}.$ (35)
We denote
$\nu=\sigma_{d}({\mathbf{\Sigma}}_{+,in})/\|{\mathbf{\Sigma}}_{+,out}\|$ (36)
and bound $\nu$ from below as follows (while explaining the details below):
$\nu\geq\frac{h({\mathbf{\Sigma}}_{+,in})\sigma_{d}({\mathbf{\Sigma}}_{in,*})}{\|{\mathbf{\Sigma}}_{+,out}\|}\geq\frac{\frac{n_{1}}{d}h({\mathbf{\Sigma}})\sigma_{d}({\mathbf{\Sigma}}_{in,*})}{{\mathcal{A}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}=\hat{\kappa}_{1}({\mathbf{\Sigma}})\frac{n_{1}\sigma_{d}({\mathbf{\Sigma}}_{in,*})}{d{\mathcal{A}}}.$ (37)
The first inequality applies (35). The second one applies Lemma 2(a) for the numerator and Lemma 2(b) for the denominator.
We use the above estimates to bound the denominator of $\bm{\hat{\kappa}_{1}}$ from above under the condition $\nu\geq 2$. We note that (37) and the assumption $\hat{\kappa}_{1}({\mathbf{\Sigma}})\geq\kappa_{1}$ (see (30)) imply that this condition is satisfied when
$\tilde{\kappa}_{1}\geq 2\frac{d{\mathcal{A}}}{n_{1}\sigma_{d}({\mathbf{\Sigma}}_{in,*})}.$ (38)
Under this assumption, we obtain the following bound, which we clarify below:
$\sigma_{1}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})\leq\sigma_{1}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})+\gamma\,\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\,\sigma_{1}([{\mathbf{P}}_{d+1,D}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})$ (39)
$=\sigma_{1}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})+\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\leq\frac{\|[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}}\|^{2}}{\sigma_{d}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}})}+\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})$
$\leq\frac{4\|{\mathbf{\Sigma}}_{+,out}\|^{2}}{\sigma_{d}({\mathbf{\Sigma}}_{+,in})-2\|{\mathbf{\Sigma}}_{+,out}\|}+\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})=\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{+,out}\|+\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+}).$
The first inequality follows from the definition of $T$ and the fact that for ${\mathbf{A}}\in S_{+}$, $\sigma_{1}({\mathbf{A}})$ is the spectral norm of ${\mathbf{A}}$. The first equality follows from the basic observation that if $\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})>0$, then $\sigma_{1}([{\mathbf{P}}_{d+1,D}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})=1$. The second inequality follows from Lemma 7 and the fact that $\Pi_{d}({\mathbf{\Sigma}}_{+})$ is of rank $d$ as follows:
$\sigma_{1}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})=\sigma_{1}\Big{(}[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}})^{-1}[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}^{\perp}}\Big{)}$
$\leq$ $\Big{\|}[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}}\Big{\|}\cdot\Big{\|}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}})^{-1}\Big{\|}\cdot\Big{\|}[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}^{\perp}}\Big{\|}=\frac{\|[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}}\|^{2}}{\sigma_{d}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}})}.$ (40)
The third inequality in (39) is verified as follows. The numerator uses the bound
$\|[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}}\|=\|[\Pi_{d}({\mathbf{\Sigma}}_{+})-{\mathbf{\Sigma}}_{+,in}]_{L_{*}^{\perp},L_{*}}\|\leq\|[\Pi_{d}({\mathbf{\Sigma}}_{+})-{\mathbf{\Sigma}}_{+,in}]\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|,$
where the last inequality is obtained by (26). The bound for the denominator applies (26) with Lemma 3(b) as follows:
$|\sigma_{d}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}})-\sigma_{d}({\mathbf{\Sigma}}_{+,in})|\leq\|[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}}-{\mathbf{\Sigma}}_{+,in}\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|.$
The last equality of (39) follows from the definition of $\nu$.
The second term in the RHS of (39) can be further bounded as follows:
$\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})=\frac{1}{D-d}\sum_{i=d+1}^{D}\sigma_{i}({\mathbf{\Sigma}}_{+})=\min_{{\mathbf{V}}\in O(D,D-d)}\operatorname{tr}({\mathbf{V}}^{T}{\mathbf{\Sigma}}_{+}{\mathbf{V}})$
$\leq\frac{1}{D-d}\operatorname{tr}({\mathbf{U}}_{L_{*}^{\perp}}^{\top}{\mathbf{\Sigma}}_{+}{\mathbf{U}}_{L_{*}^{\perp}})=\frac{1}{D-d}\operatorname{tr}({\mathbf{U}}_{L_{*}^{\perp}}^{\top}{\mathbf{\Sigma}}_{+,out}{\mathbf{U}}_{L_{*}^{\perp}})\leq\frac{n_{0}}{D-d}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}).$ (41)
The second equality is the second extremal partial trace formula in [9, Proposition 1.3.4] and the second inequality applies the trace bound in (18). The third equality is because the inliers lie on $L_{*}$ and thus ${\mathbf{U}}_{L_{*}^{\perp}}^{\top}{\mathbf{\Sigma}}_{+,in}{\mathbf{U}}_{L_{*}^{\perp}}={\mathbf{0}}$.
Combining the bound on ${\mathbf{\Sigma}}_{+,out}$ in (18), (39) and (3.1.4), we conclude the upper bound on the denominator in (33):
$\sigma_{1}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})\leq\frac{4}{\nu-2}{\mathcal{A}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})+\gamma\frac{n_{0}}{D-d}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}).$ (42)
Estimate of the numerator of $\bm{\hat{\kappa}_{1}}$: We note that
$\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}}-[g_{1}(T({\mathbf{\Sigma}}))]_{L_{*},L_{*}})\|=\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}^{\perp}}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})^{-1}[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}}\|\leq\frac{\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}^{\perp}}\|^{2}}{\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})}.$ (43)
Indeed, the first equality follows by the definition of $g_{1}$ and the first inequality uses the sub-multiplicative property of the spectral norm and basic estimates, similar to verifying the second line of (3.1.4).
In addition, we note that
$\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}}-[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}\|=\|[T({\mathbf{\Sigma}})-{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}\|\leq\|T({\mathbf{\Sigma}})-{\mathbf{\Sigma}}_{+,in}\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|,$ (44)
where the last inequality applies (25). Next, we bound $h(T({\mathbf{\Sigma}}))$ using arguments detailed below:
$h(T({\mathbf{\Sigma}}))$ $=\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}[g_{1}(T({\mathbf{\Sigma}}))]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5})$
$\geq\sigma_{d}({\mathbf{\Sigma}}_{in,*}^{-0.5}[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}{\mathbf{\Sigma}}_{in,*}^{-0.5})-\big{\|}{\mathbf{\Sigma}}_{in,*}^{-0.5}([{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}-[g_{1}(T({\mathbf{\Sigma}}))]_{L_{*},L_{*}}){\mathbf{\Sigma}}_{in,*}^{-0.5}\big{\|}$
$\geq h({\mathbf{\Sigma}}_{+,in})-\frac{\|[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}-[g_{1}(T({\mathbf{\Sigma}}))]_{L_{*},L_{*}}\|}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}\geq h({\mathbf{\Sigma}}_{+,in})-\frac{2\|{\mathbf{\Sigma}}_{+,out}\|+\frac{\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}^{\perp}}\|^{2}}{\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})}}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}.$ (45)
The first inequality follows from Lemma 3(b). The second inequality first applies the definition of $h$ and then to the second term the following argument, which is similar to the one in (35):
$\|{\mathbf{\Sigma}}_{in,*}^{-0.5}{\mathbf{X}}{\mathbf{\Sigma}}_{in,*}^{-0.5}\|\leq\|{\mathbf{\Sigma}}_{in,*}^{-0.5}\|\cdot\|{\mathbf{X}}\|\cdot\|{\mathbf{\Sigma}}_{in,*}^{-0.5}\|=\frac{\|{\mathbf{X}}\|}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}.$
The last inequality applies the triangle inequality with (43) and (44).
The component $\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})$ in the RHS of (3.1.4) as follows:
$\sigma_{D-d}\big{(}[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\big{)}\geq\sigma_{D}\big{(}T({\mathbf{\Sigma}})\big{)}=\sigma_{D}\big{(}T_{2}({\mathbf{\Sigma}}_{+},\gamma)\big{)}=\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\geq\gamma\sigma_{D}({\mathbf{\Sigma}})\,{\mathcal{S}},$ (46)
where the first inequality follows from Lemma 3(a) and the second inequality follows from Lemma 2(c). In addition, we bound another component of the RHS of (3.1.4) using the triangle inequality and (25) as follows:
$\|[{T}({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\|\leq 2\|{\mathbf{\Sigma}}_{+,out}\|+\|[{\mathbf{\Sigma}}_{+,in}]_{L_{*}^{\perp},L_{*}^{\perp}}\|=2\|{\mathbf{\Sigma}}_{+,out}\|.$
Combining the above two equations and the bound of $\|{\mathbf{\Sigma}}_{+,out}\|$ in (18) yields
$\frac{\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}^{\perp}}\|^{2}}{\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})}\leq\frac{(2\|{\mathbf{\Sigma}}_{+,out}\|)^{2}}{\gamma\sigma_{D}({\mathbf{\Sigma}}){\mathcal{S}}}=\frac{4\|{\mathbf{\Sigma}}_{+,out}\|{\mathcal{A}}}{\gamma{\mathcal{S}}}\cdot\frac{\sigma_{1}([{\mathbf{\Sigma}}]_{L_{*}^{\perp},L_{*}^{\perp}})}{\sigma_{D}({\mathbf{\Sigma}})}=\frac{4\|{\mathbf{\Sigma}}_{+,out}\|{\mathcal{A}}}{\gamma{\mathcal{S}}}\,\hat{\kappa}_{2}({\mathbf{\Sigma}}).$ (47)
Plugging the above bound in (3.1.4) results in the following bound on the numerator of (33):
$h(T({\mathbf{\Sigma}}))\geq h({\mathbf{\Sigma}}_{+,in})-\frac{2\|{\mathbf{\Sigma}}_{+,out}\|+\frac{4\|{\mathbf{\Sigma}}_{+,out}\|\,{\mathcal{A}}\,\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}}}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}=h({\mathbf{\Sigma}}_{+,in})-\Big{(}{2+\frac{4\,{\mathcal{A}}\,\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}}}\Big{)}\frac{\|{\mathbf{\Sigma}}_{+,out}\|}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}.$ (48)
Conclusion of the proof that $\bm{\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq C_{0}\hat{\kappa}_{1}({\mathbf{\Sigma}})}$: Combining (42) and (48), we obtain the following bound when $\nu>2$:
$\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq\frac{h({\mathbf{\Sigma}}_{+,in})-\Big{(}{2+\frac{4{\mathcal{A}}\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}}}\Big{)}\frac{\|{\mathbf{\Sigma}}_{+,out}\|}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}}{\frac{4}{\nu-2}{\mathcal{A}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})+\gamma\frac{n_{0}}{D-d}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}$
$=$ $\frac{h({\mathbf{\Sigma}}_{+,in})-\Big{(}{2+\frac{4{\mathcal{A}}\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}}}\Big{)}\frac{{\sigma_{d}({\mathbf{\Sigma}}_{+,in})}}{\nu\sigma_{d}({\mathbf{\Sigma}}_{in,*})}}{\frac{4}{\nu-2}{\mathcal{A}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})+\gamma\frac{n_{0}}{D-d}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}\geq\frac{h({\mathbf{\Sigma}}_{+,in})-(2+\frac{4{\mathcal{A}}\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}})\frac{\kappa_{in}h({\mathbf{\Sigma}}_{+,in})}{\nu}}{\frac{4}{\nu-2}{\mathcal{A}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})+\gamma\frac{n_{0}}{D-d}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}$
$=$ $\frac{1-(2+\frac{4{\mathcal{A}}\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}})\frac{\kappa_{in}}{\nu}}{\frac{4}{\nu-2}{\mathcal{A}}+\gamma\frac{n_{0}}{D-d}}\frac{h({\mathbf{\Sigma}}_{+,in})}{\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}\geq\frac{1-(2+\frac{4{\mathcal{A}}\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}})\frac{\kappa_{in}}{\nu}}{\frac{4}{\nu-2}\,{\mathcal{A}}+\gamma\frac{n_{0}}{D-d}}\frac{n_{1}}{d}\hat{\kappa}_{1}({\mathbf{\Sigma}}).$ (49)
The first equality uses the definition of $\nu$ in the last multiplicative term of the numerator. The second inequality applies (34), that is, $\sigma_{d}({\mathbf{\Sigma}}_{+,in})\leq h({\mathbf{\Sigma}}_{+,in})\sigma_{1}({\mathbf{\Sigma}}_{in,*})$, and the definition of $\kappa_{in}$, that is, $\kappa_{in}=\sigma_{1}({\mathbf{\Sigma}}_{in,*})/\sigma_{d}({\mathbf{\Sigma}}_{in,*})$. The last inequality follows from Lemma 2(a) and the definition $\hat{\kappa}_{1}({\mathbf{\Sigma}})=h({\mathbf{\Sigma}})/{\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})}$. When $\kappa_{1}$ satisfies the lower bound in (3) with the constant $C$, the requirement on $\nu$ in (37) becomes
$\nu\geq C\left(\kappa_{in}+\frac{{\mathcal{A}}}{\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}}+\frac{\kappa_{2}{\mathcal{A}}}{\gamma{\mathcal{S}}}(1+\kappa_{in})\right).$ (50)
Since $C$ can be chosen sufficiently large and since $\kappa_{in}\geq 1$, $\nu$ can be chosen sufficiently large and satisfy the assumption used earlier that $\nu>2$. We even assume here that $\nu\geq 4$ (which holds when $C\geq 4$) and obtain the estimate
$\frac{4}{\nu-2}{\mathcal{A}}\leq\frac{8}{\nu}{\mathcal{A}}\leq\frac{8}{C}\Big{(}\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}\Big{)}\ \text{ and }\ \Big{(}2+\frac{4{\mathcal{A}}\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\gamma{\mathcal{S}}}\Big{)}\frac{\kappa_{in}}{\nu}\leq\frac{2}{C}+\frac{4}{C}=\frac{6}{C}.$
The estimates in the above equation combined with (49) lead to the bound
$\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq\frac{1-(\frac{6}{C})}{\frac{8}{C}\Big{(}\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}\Big{)}+\gamma\frac{n_{0}}{D-d}}\frac{n_{1}}{d}\hat{\kappa}_{1}({\mathbf{\Sigma}}).$ (51)
To conclude the proof, we show that when $C$ is sufficiently large,
$\frac{1-\frac{6}{C}}{\frac{8}{C}\Big{(}\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}\Big{)}+\gamma\frac{n_{0}}{D-d}}\frac{n_{1}}{d}>C_{0}.$ (52)
To show this, we simplify (52) as follows. Its LHS becomes
$\frac{1-\frac{6}{C}}{\frac{8}{C}\Big{(}\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}\Big{)}}\frac{n_{1}}{d}=\frac{1-\frac{6}{C}}{\frac{8}{C}(\frac{\text{DS-SNR}}{\gamma}-1)+1}\frac{\text{DS-SNR}}{\gamma}$
and in view of (29) its RHS is $2\cdot\frac{\text{DS-SNR}}{\text{DS-SNR}+\gamma}$. We thus note that (52) holds when
$\frac{1-\frac{6}{C}}{\frac{8}{C}(\frac{\text{DS-SNR}}{\gamma}-1)+1}\frac{\text{DS-SNR}}{\gamma}>\frac{2\cdot\text{DS-SNR}}{\text{DS-SNR}+\gamma},$
that is,
$\frac{6}{C}\left(\frac{\text{DS-SNR}}{\gamma}+1\right)+\frac{16}{C}\left(\frac{\text{DS-SNR}}{\gamma}-1\right)<\frac{\text{DS-SNR}}{\gamma}-1,$
or in other words
$C>\frac{22\cdot\text{DS-SNR}-10\cdot\gamma}{\text{DS-SNR}-\gamma}.$ (53)
Therefore, there exists $C$ sufficiently large (specified in (53)) such that $\hat{\kappa}_{1}(T({\mathbf{\Sigma}}))\geq C_{0}\hat{\kappa}_{1}({\mathbf{\Sigma}})$.
Proof that $\bm{\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))\geq\max(\hat{\kappa}_{2}({\mathbf{\Sigma}})},7)$: Recall that
$\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))=\frac{\sigma_{1}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})}{\sigma_{D}(T({\mathbf{\Sigma}}))}.$
We note that (39) bounds the numerator of $\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))$. Next, we bound the denominator of $\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))$, where the main tool is Lemma 5 and at last we use these estimates to conclude that ${\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))\leq\max(\hat{\kappa}_{2}({\mathbf{\Sigma}})},7)$.
Estimate of the denominator of $\bm{\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))}$: We derive (57) below and then apply Lemma 5. We first note that if
$C>58,$ (54)
then
$(\nu-2)\gamma{\mathcal{S}}=\frac{\nu-2}{\nu}\nu\gamma{\mathcal{S}}\geq\frac{\nu-2}{\nu}2C\kappa_{2}{\mathcal{A}}\geq\frac{C-2}{C}\cdot 2C\kappa_{2}{\mathcal{A}}=2(C-2){\mathcal{A}}\kappa_{2}\geq\frac{2(C-2)}{7}{\mathcal{A}}\max(\kappa_{2},7)\geq 16{\mathcal{A}}\hat{\kappa}_{2}({\mathbf{\Sigma}}).$ (55)
The first inequality follows from (50) (using only the third term in the sum as a lower bound of $\nu$) and the fact $\kappa_{in}\geq 1$. The second inequality follows from the fact that $\nu\geq C$, which is a consequence of (50) (using only the first term in the sum as a lower bound for $\nu$) and the fact $\kappa_{in}\geq 1$. The last inequality follows from $C\geq 58$ and the condition $\hat{\kappa}_{2}({\mathbf{\Sigma}})\leq\max(\kappa_{2},7)$ in (30).
Next, we verify the following bound:
$(2\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}^{\perp}}\|)^{2}$ $\leq 16(\|{\mathbf{\Sigma}}_{+,out}\|^{2})=16\|{\mathbf{\Sigma}}_{+,out}\|\cdot\|{\mathbf{\Sigma}}_{+,out}\|\leq 8{\mathcal{A}}\sigma_{1}({\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}})\cdot\|{\mathbf{\Sigma}}_{+,out}\|$
$\leq\frac{(\nu-2)\gamma{\mathcal{S}}}{\hat{\kappa}_{2}({\mathbf{\Sigma}})}{\sigma_{1}({\mathbf{\Sigma}}_{L_{+}^{\perp},L_{+}^{\perp}})}\cdot\|{\mathbf{\Sigma}}_{+,out}\|=(\nu-2)\gamma\sigma_{D}({\mathbf{\Sigma}}){\mathcal{S}}\|{\mathbf{\Sigma}}_{+,out}\|=\Big{(}(\nu-2)\|{\mathbf{\Sigma}}_{+,out}\|\Big{)}\Big{(}\gamma\sigma_{D}({\mathbf{\Sigma}}){\mathcal{S}}\Big{)}.$ (56)
The first inequality follows from (25) and the fact that $[{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}^{\perp}}$ is a zero matrix. The second inequality applies (18). The third inequality applies (55). The next equality applies the definition of $\hat{\kappa}_{2}({\mathbf{\Sigma}})$.
We bound the first multiplicative term in the RHS of (56) as follows:
$(\nu-2)\|{\mathbf{\Sigma}}_{+,out}\|=\sigma_{d}([{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}}-2\|{\mathbf{\Sigma}}_{+,out}\|\leq\sigma_{d}([T({\mathbf{\Sigma}})]_{L_{*},L_{*}}).$
The second multiplicative term in the RHS of (56) was already bounded in (46) as follows:
$\gamma\sigma_{D}({\mathbf{\Sigma}})\,{\mathcal{S}}\leq\sigma_{D-d}\big{(}[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\big{)}.$
The above three equations thus result in the following bound:
$\|[T({\mathbf{\Sigma}})]_{L_{*},L_{*}^{\perp}}\|\leq\sqrt{\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})\sigma_{d}([T({\mathbf{\Sigma}})]_{L_{*},L_{*}})}/2.$ (57)
The above equation is the condition stated in (23) of Lemma 5 for $T({\mathbf{\Sigma}})$. This lemma thus implies that
$\sigma_{D}(T({\mathbf{\Sigma}}))\geq\frac{1}{3}\min\Big{(}\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}),\sigma_{d}([T({\mathbf{\Sigma}})]_{L_{*},L_{*}})\Big{)}.$ (58)
Conclusion of the Proof that $\bm{\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))\geq\max(\hat{\kappa}_{2}({\mathbf{\Sigma}})},7)$: We first note that (58) implies
$\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))=\frac{\|[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\|}{\sigma_{D}(T({\mathbf{\Sigma}}))}\leq 3\max\left(\frac{\|[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\|}{\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})}\,,\,\frac{\|[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\|}{\sigma_{d}([T({\mathbf{\Sigma}})]_{L_{*},L_{*}})}\right).$ (59)
We bound the second argument of the above maximum as follows:
$\frac{\|[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\|}{\sigma_{d}([T({\mathbf{\Sigma}})]_{L_{*},L_{*}})}\leq\frac{\|[{\mathbf{\Sigma}}_{+}]_{L_{*}^{\perp},L_{*}^{\perp}}\|+2\|{\mathbf{\Sigma}}_{+,out}\|}{\sigma_{d}([{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}})-2\|{\mathbf{\Sigma}}_{+,out}\|}\leq\frac{3\|{\mathbf{\Sigma}}_{+,out}\|}{\sigma_{d}([{\mathbf{\Sigma}}_{+,in}]_{L_{*},L_{*}})-2\|{\mathbf{\Sigma}}_{+,out}\|}=\frac{3}{\nu-2}.$ (60)
For the numerator in the first inequality we used the triangle inequality and Lemma 6. For the denominator in the second inequality we used Lemma 3(b) and also the triangle inequality and Lemma 6. We can divide the two bounds as long as the denominator is positive, which holds when $\nu>2$ (recall the definition of $\nu$ in (36)). The second inequality uses the fact $[{\mathbf{\Sigma}}_{+,in}]_{L_{*}^{\perp},L_{*}^{\perp}}={\mathbf{0}}$ and the last equality uses the definition of $\nu$.
To bound the first argument of the maximum in (59), we first note that
$\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})$ $\geq\sigma_{D-d}([T({\mathbf{\Sigma}})-\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})-\sigma_{1}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})=\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})-\sigma_{1}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}^{\perp}})$
$\geq$ $\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})-\frac{\|[\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*}^{\perp},L_{*}}\|^{2}}{\sigma_{d}([\Pi_{d}({\mathbf{\Sigma}}_{+})]_{L_{*},L_{*}})}\geq\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})-\frac{4\|{\mathbf{\Sigma}}_{+,out}\|^{2}}{\sigma_{d}({\mathbf{\Sigma}}_{+,in})-2\|{\mathbf{\Sigma}}_{+,out}\|}$
$=$ $\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})-\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{+,out}\|.$
The first inequality follows from Lemma 3(b). The rest of the steps follow arguments similar to deriving (39). Using this inequality and the bound on $\|[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\|\equiv\sigma_{1}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})$ in (39), we bound the first argument of the maximum above:
$\frac{\|[T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}}\|}{\sigma_{D-d}([T({\mathbf{\Sigma}})]_{L_{*}^{\perp},L_{*}^{\perp}})}\leq\frac{\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{+,out}\|+\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})}{-\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{+,out}\|+\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})}.$
To further bound the RHS of the above inequality, we note that for $a=\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{+,out}\|$ and $b=\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})$, $(a+b)/(-a+b)\equiv 1+2((b/a)-1)^{-1}$ can be bounded using a lower bound of $b$ and an upper bound of $a$ and verifying that the ratio, while making sure that the lower bound for $b$ is larger than the upper bound of $a$. In order to bound $b$ from below, we apply (19) and the definition of ${\mathcal{S}}$:
$\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})\geq\sigma_{d+1,D}\Big{(}\sigma_{D}({\mathbf{\Sigma}})\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}\Big{)}=\sigma_{D}({\mathbf{\Sigma}})\sigma_{d+1,D}\Big{(}\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}\mathbf{x}^{\top}}{\|\mathbf{x}\|^{2}}\Big{)}=\sigma_{D}({\mathbf{\Sigma}}){\mathcal{S}}.$
To upper bound $a$, we apply the following bound of Lemma 2(b):
$\|{\mathbf{\Sigma}}_{+,out}\|\leq\|{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\|\,{\mathcal{A}}.$
Finally, to make sure that our bound is valid so At last we make sure that the expression $(a+b)/(-a+b)$ is positive after using these bounds, so our argument is valid. We assume that $C>28$ and use the bounds $\nu>C$ and $\nu>2C\frac{\kappa_{2}{\mathcal{A}}}{\gamma{\mathcal{S}}}$ (already explained when bounding (55)) and the assumption $\hat{\kappa}_{2}({\mathbf{\Sigma}})\leq\max(\kappa_{2},7)$ in (30) to obtain
$\frac{4}{\nu-2}\hat{\kappa}_{2}({\mathbf{\Sigma}}){\mathcal{A}}\leq\frac{8}{\nu}\hat{\kappa}_{2}({\mathbf{\Sigma}}){\mathcal{A}}\leq\frac{8}{\nu}\max(\kappa_{2},7){\mathcal{A}}\leq\frac{28}{C}\gamma{\mathcal{S}}.$
Applying this bound and the above estimates, we conclude the last requirement on $a$ and $b$:
above lower bound of $\frac{\text{above lower bound of }\gamma\sigma_{d+1,D}({\mathbf{\Sigma}}_{+})}{\text{upper bound of }\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{+,out}\|}=\frac{\gamma\sigma_{D}({\mathbf{\Sigma}}){\mathcal{S}}}{\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\|{\mathcal{A}}}=\frac{\gamma{\mathcal{S}}}{\frac{4}{\nu-2}\hat{\kappa}_{2}({\mathbf{\Sigma}}){\mathcal{A}}}\geq\frac{C}{28}\geq 1.$
All above estimates (staring at (59)) yield
$\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))\leq 3\,\max\left(\frac{\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\|{\mathcal{A}}+\gamma\sigma_{D}({\mathbf{\Sigma}}){\mathcal{S}}}{-\frac{4}{\nu-2}\|{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\|{\mathcal{A}}+\gamma\sigma_{D}({\mathbf{\Sigma}}){\mathcal{S}}},\frac{3}{\nu-2}\right)=3\,\max\left(\frac{\frac{4}{\nu-2}\hat{\kappa}_{2}({\mathbf{\Sigma}}){\mathcal{A}}+\gamma{\mathcal{S}}}{-\frac{4}{\nu-2}\hat{\kappa}_{2}({\mathbf{\Sigma}}){\mathcal{A}}+\gamma{\mathcal{S}}},\frac{3}{\nu-2}\right),$ (61)
and the RHS of (61) is bounded by
$3\max\Big{(}\frac{28/C+1}{-28/C+1},\frac{3}{\nu-2}\Big{)}\leq 3\max\Big{(}\frac{28/C+1}{-28/C+1},\frac{3}{C-2}\Big{)}$
which is bounded by $7$ for a sufficiently large $C$. In fact, combining the bounds of $C$ in (53) and (54), it is sufficient to choose the final value of $C$ such that
$C\geq\max\Big{(}58,\frac{22\cdot\text{DS-SNR}-10\cdot\gamma}{\text{DS-SNR}-\gamma}\Big{)}.$ (62)
As a result, $\hat{\kappa}_{2}(T({\mathbf{\Sigma}}))\leq 7\leq\max(7,\kappa_{2})$.
3.2 Proof of Theorem 2
We denote by TME+STE the STE algorithm initialized by TME. We will prove the following three claims that will imply the statement of Theorem 2.
  • TME+STE recovers $L_{*}$ under a general model where the inliers lie on the $d$-subspace $L_{*}$, both ${\mathcal{X}}_{in}$ and $\{{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}\,:\,\mathbf{x}\in{\mathcal{X}}_{out}\}$ lie in general position, that is, any $k$-dimensional subspace contains at most $k$ points, and either ($N\geq D$ and DS-SNR$>1$) or ($N>D$ and DS-SNR$=1$).
  • TME does not recover $L_{*}$ for data generated by a generalized haystack model, where ${\mathbf{\Sigma}}^{(out)}_{L_{*},L_{*}^{\perp}}\neq 0$ and DS-SNR$<1$.
  • TME+STE recovers $L_{*}$ when $\eta\leq$DS-SNR$<1$ for data generated by a generalized haystack model.
Note that the first statement is rather general and clearly holds for the special case of data sampled by the generalized haystack model. We also remark that our generalized haystack is not as general as the one of [8]. Nonetheless, we still call it generalized since the covariance of the outliers is not symmetric. The case of symmetric covariance for outliers is commonly assumed in the haystack model, which makes it very limited. For this simpler common model, the condition in the second statement above will not hold and TME will still be able to asymptotically recover the subspace.
Below we verify in order these three claims.
TME+STE recovers $L_{*}$ when DS-SNR1absent1$\geq$ 1≥ 1: If DS-SNR$>1$, [13] showed that the range of the TME solution, ${\mathbf{\Sigma}}^{(TME)}$, is $L_{*}$ and thus TME recovers $L_{*}$. We note that $T_{1}({\mathbf{\Sigma}}^{(TME)})={\mathbf{\Sigma}}^{(TME)}$ and since $\operatorname{rank}({\mathbf{\Sigma}}^{(TME)})=d$ then also $T_{2}({\mathbf{\Sigma}}^{(TME)},\gamma)={\mathbf{\Sigma}}^{(TME)}$. Consequently, $T({\mathbf{\Sigma}}^{(TME)})=T_{2}(T_{1}({\mathbf{\Sigma}}^{(TME)}),\gamma)={\mathbf{\Sigma}}^{(TME)}$. That is, if for STE ${\mathbf{\Sigma}}^{(0)}={\mathbf{\Sigma}}^{(TME)}$, then ${\mathbf{\Sigma}}^{(k)}={\mathbf{\Sigma}}^{(TME)}$ for all $k\geq 1$ and thus TME+STE also recovers $L_{*}$.
If DS-SNR$=1$, the following argument shows that the range of the TME solution is also $L_{*}$, and thus implies that TME and TME+STE recover $L_{*}$. For simplicity, we denote the TME solution by ${\mathbf{\Sigma}}$ only and note that as stated in (20) it satisfies:
${\mathbf{\Sigma}}=\frac{D}{N}\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}{\mathbf{\Sigma}}^{-1}\mathbf{x}}.$
We assume first that ${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}$ is nonsingular. We derive the following matrix inequality using the above formula, but replacing the sum over $\mathbf{x}\in{\mathcal{X}}$ to $\mathbf{x}\in{\mathcal{X}}_{out}$ since ${\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}\mathbf{x}^{T}{\mathbf{U}}_{L_{*}^{\perp}}=0$ for $\mathbf{x}\in{\mathcal{X}}_{in}$, and the same arguments used to obtain the top line of (21):
${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}=\frac{D}{N}\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}\mathbf{x}^{T}{\mathbf{U}}_{L_{*}^{\perp}}}{\mathbf{x}^{T}({\mathbf{\Sigma}})^{-1}\mathbf{x}}\preceq\frac{D}{N}\lim_{t\to\infty}\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}\mathbf{x}^{T}{\mathbf{U}}_{L_{*}^{\perp}}}{\mathbf{x}^{T}({\mathbf{\Sigma}}+t\,{\mathbf{P}}_{L_{*}})^{-1}\mathbf{x}}=\frac{D}{N}\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}\mathbf{x}^{T}{\mathbf{U}}_{L_{*}^{\perp}}}{({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})^{T}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-1}({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})}.$ (63)
Since $\gamma=1$, $D/N=(D-d)/n_{0}$ and (63) implies
${\mathbf{I}}_{(D-d)\times(D-d)}\preceq\frac{D}{N}\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\mathbf{x}^{T}{\mathbf{U}}_{L_{*}^{\perp}}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-0.5}}{({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})^{T}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-1}({\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x})}=\frac{D-d}{n_{0}}\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\Big{(}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\Big{)}\Big{(}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\Big{)}^{T}}{\Big{(}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\Big{)}^{T}\Big{(}{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}\Big{)}}.$ (64)
Since the traces of both the LHS and RHS in (64) are $D-d$, equality must hold in (64) and consequently also in (63). Note that equality in (63) suggests that $\lim_{t\to\infty}\mathbf{x}^{T}\big{(}{\mathbf{\Sigma}}^{-1}-({\mathbf{\Sigma}}+t\,{\mathbf{P}}_{L_{*}})^{-1}\big{)}\mathbf{x}=0$ for all $\mathbf{x}\in{\mathcal{X}}_{out}$. Since ${\mathbf{\Sigma}}^{-1}\succeq\lim_{t\to\infty}({\mathbf{\Sigma}}+t\,{\mathbf{P}}_{L_{*}})^{-1}$, the equality holds only when every $\mathbf{x}\in{\mathcal{X}}_{out}$ lies in the kernel of $\lim_{t\to\infty}\left({\mathbf{\Sigma}}^{-1}-({\mathbf{\Sigma}}+t{\mathbf{P}}_{L_{*}})^{-1}\right)$, which is a violation of the general position assumption that ${\mathcal{X}}_{out}$ is not contained in a subspace within ${\mathbb{R}}^{D}$.
We thus conclude that ${\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}$ is singular and so is ${\mathbf{\Sigma}}$. We next prove the following property.
Lemma 9.
Let $L_{0}$ denote the range of ${\mathbf{\Sigma}}$. Then, $L_{0}\subseteq L_{*}$.
Proof of Lemma 9.
We note that the TME iteration at step $k$, ${\mathbf{\Sigma}}^{(k),TME}$, satisfies
$g_{1}({\mathbf{\Sigma}}^{(k+1),TME})=g_{1}\Big{(}\frac{D}{N}\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}({\mathbf{\Sigma}}^{(k),TME})^{-1}\mathbf{x}}\Big{)}\succeq\frac{D}{N}\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}({\mathbf{\Sigma}}^{(k),TME})^{-1}\mathbf{x}}$
$\succeq$ $\frac{D}{N}\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}(g_{1}({\mathbf{\Sigma}}^{(k),TME}))^{-1}\mathbf{x}}=\frac{d}{n_{1}}\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}(g_{1}({\mathbf{\Sigma}}^{(k),TME}))^{-1}\mathbf{x}}.$ (65)
where the first inequality follows from Lemma 1(b) and the fact that
$\sum_{\mathbf{x}\in{\mathcal{X}}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}({\mathbf{\Sigma}}^{(k),TME})^{-1}\mathbf{x}}-\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}({\mathbf{\Sigma}}^{(k),TME})^{-1}\mathbf{x}}=\sum_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}({\mathbf{\Sigma}}^{(k),TME})^{-1}\mathbf{x}}\in S_{+}.$
The second inequality follows from ${\mathbf{\Sigma}}^{(k),TME}\succeq g_{1}({\mathbf{\Sigma}}^{(k),TME})$ (since ${\mathbf{\Sigma}}-g_{1}({\mathbf{\Sigma}})\equiv g_{2}({\mathbf{\Sigma}})\succeq{\mathbf{0}}$ by Lemma 1).
To simplify our arguments we recall the following property of the TME solution [10]: If $\tilde{{\mathbf{\Sigma}}}^{(TME)}$ and ${{\mathbf{\Sigma}}}^{(TME)}$ are the TME solutions to the sets $\tilde{{\mathbf{X}}}=\{{\mathbf{A}}\mathbf{x}:\mathbf{x}\in{\mathcal{X}}\}$ and ${\mathcal{X}}$, then
$\tilde{{\mathbf{\Sigma}}}^{(TME)}={\mathbf{A}}{{\mathbf{\Sigma}}}^{(TME)}{\mathbf{A}}^{T}.$ (66)
WLOG we may assume that the TME solution for $\{{\mathbf{U}}_{L_{*}^{\perp}}\mathbf{x}:\mathbf{x}\in{\mathcal{X}}_{out}\}$ is the identity; otherwise we can apply the linear transformation ${\mathbf{A}}:={{{\mathbf{\Sigma}}}^{(TME)}}^{-0.5}$ such that the new solution satisfies $\tilde{{\mathbf{\Sigma}}}^{(TME)}={\mathbf{A}}{{\mathbf{\Sigma}}}^{(TME)}{\mathbf{A}}^{T}={\mathbf{I}}$. Then (20) and the above assumption imply
$\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}^{T}\mathbf{x}}{\|\mathbf{x}\|^{2}}=\frac{n_{1}}{d}{\mathbf{P}}_{L_{*}}.$ (67)
We note that
$\sigma_{d}(g_{1}({\mathbf{\Sigma}}^{(k+1),TME}))\geq\sigma_{d}\Big{(}\frac{d}{n_{1}}\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}(g_{1}({\mathbf{\Sigma}}^{(k),TME}))^{-1}\mathbf{x}}\Big{)}\geq\sigma_{d}\Big{(}\frac{d}{n_{1}}\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}\mathbf{x}^{T}}{\mathbf{x}^{T}(\sigma_{d}^{-1}(g_{1}({\mathbf{\Sigma}}^{(k),TME})))\mathbf{x}}\Big{)}$
$=$ $\sigma_{d}(g_{1}({\mathbf{\Sigma}}^{(k),TME}))\sigma_{d}\Big{(}\frac{d}{n_{1}}\sum_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\mathbf{x}^{T}\mathbf{x}}{\|\mathbf{x}\|^{2}}\Big{)}=\sigma_{d}(g_{1}({\mathbf{\Sigma}}^{(k),TME})),$
where the first inequality uses (65) and the last equality uses (67).
Therefore $\{\sigma_{d}(g_{1}({\mathbf{\Sigma}}^{(k),TME}))\}_{k\geq 0}$ is a nondecreasing sequence and thus $g_{1}({\mathbf{\Sigma}})=\lim_{k\rightarrow\infty}g_{1}({\mathbf{\Sigma}}^{(k),TME})$ is positive definite. Moreover, any nonzero vector ${\mathbf{u}}$ that is not in $L_{*}^{\perp}$ satisfies ${\mathbf{u}}^{T}g_{1}({\mathbf{\Sigma}}){\mathbf{u}}>0$. We recall that Lemma 1 implies that for ${\mathbf{\Sigma}}\in S_{+}$, ${\mathbf{\Sigma}}-g_{1}({\mathbf{\Sigma}})=g_{2}({\mathbf{\Sigma}})\succeq{\mathbf{0}}$. Combining this observation with the latter one yields ${\mathbf{u}}^{T}{\mathbf{\Sigma}}{\mathbf{u}}>0$. Consequently, the kernel of ${\mathbf{\Sigma}}$ must lie in $L_{*}^{\perp}$ and its range $L_{0}$ contains $L_{*}$. ∎
Then TME over ${\mathcal{X}}$ is equivalent to TME over the set ${\mathcal{X}}\cap L_{0}$, as points outside $L_{0}$ should have weight $01/\mathbf{x}^{T}{\mathbf{\Sigma}}^{-1}\mathbf{x}=$ 0. In particular, we have $D=\dim(L_{0})$. By the general position assumption, $|{\mathcal{X}}\cap L_{0}|\leq n_{1}+\dim(L_{0})-d$, and as a result, the DS-SNR for $L_{*}$ is larger than $1$ and [13] showed in this case that the range of the TME solution is $L_{*}$.
Lastly, we remark that in our setting with $N>D$ there is no other subspace than $L_{*}$ with more than $d$ points on it (this will contradict the assumptions of general position and $N>D$). However, if $N=D$ and $DS-SNR=1$, we may have at least two $d$-subspaces with more than $d$ points. Our statement thus avoids this case.
TME does not recover $L_{*}$ if ${\mathbf{\Sigma}}^{(out)}_{L_{*},L_{*}^{\perp}}\neq 0$ and DS-SNR$<1$
We first state a lemma, which clarifies properties of the TME solution under the generalized haystack model.
Lemma 10.
For data generated by the generalized haystack model and $N\to\infty$, whereas $d$, $D$, $n_{1}/N$ and $n_{0}/N$ are fixed, the following two properties hold:
(a) If ${\mathbf{\Sigma}}^{(out)}_{L_{*},L_{*}^{\perp}}$=0, the TME solution ${\mathbf{\Sigma}}^{(TME)}$ satisfies
${\mathbf{\Sigma}}^{(TME)}_{L_{*},L_{*}^{\perp}}={\mathbf{0}}\ \text{ and }\ {\mathbf{\Sigma}}^{(TME)}_{L_{*}^{\perp},L_{*}}={\mathbf{0}}.$
(b) The TME solution, with a particular choice of scale factor, satisfies the following conditions:
[…]
where $C_{1}$ and $C_{2}$ are constants depending on the condition numbers of ${\mathbf{\Sigma}}^{(out)}$ and ${\mathbf{\Sigma}}^{(in)}$.
We prove this lemma in §3.2.1. Here we use this lemma to conclude the above statement. For
${\mathbf{A}}=\begin{pmatrix}{\mathbf{I}}&-{\mathbf{\Sigma}}^{(out)}_{L_{*},L_{*}^{\perp}}({\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}})^{-1}\\ 0&{\mathbf{I}}\end{pmatrix},$
the set $\tilde{{\mathcal{X}}}=\{{\mathbf{A}}\mathbf{x}:\mathbf{x}\in{\mathcal{X}}\}$ has an “outlying covariance” ${\mathbf{A}}{\mathbf{\Sigma}}^{(out)}{\mathbf{A}}^{T}$, which has a block diagonal structure, and “inlying covariance” ${\mathbf{A}}{\mathbf{\Sigma}}^{(in)}{\mathbf{A}}^{T}={\mathbf{\Sigma}}^{(in)}$. Applying Lemma 10(a), the ${\mathbf{A}}$-transformed TME solution, $\tilde{{\mathbf{\Sigma}}}^{TME}$, satisfies $\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*},L_{*}^{\perp}}=0$.
By (66), ${{\mathbf{\Sigma}}}^{TME}={\mathbf{A}}\tilde{{\mathbf{\Sigma}}}^{TME}{\mathbf{A}}^{T}$, and it does not have the same block diagonal structure when ${\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}}\neq 0$. As a result, its top $d$ eigenvectors do not span $L_{*}.$
TME+STE recovers $L_{*}$ when $\eta\leq$DS-SNR$<1$. Note that the operator ${\mathbf{A}}$ satisfies
$\|{\mathbf{A}}\|\leq 2+\|{\mathbf{\Sigma}}^{(out)}_{L_{*},L_{*}^{\perp}}({\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}})^{-1}\|\leq 2+\kappa_{out}$
and
$\|{\mathbf{A}}^{-1}\|=\left\|\begin{pmatrix}{\mathbf{I}}&{\mathbf{\Sigma}}^{(out)}_{L_{*},L_{*}^{\perp}}({\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}})^{-1}\\ 0&{\mathbf{I}}\end{pmatrix}\right\|\leq 2+\kappa_{out}.$
Applying Lemma 10(b) to the ${\mathbf{A}}$-transformed solution, $\tilde{{\mathbf{\Sigma}}}^{TME}$, results in
$\frac{\sigma_{d}(\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*},L_{*}})}{\sigma_{1}(\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})}\geq O\left(\frac{1}{1-\text{DS-SNR}}\right).$ (68)
The application of ${{\mathbf{\Sigma}}}^{TME}={\mathbf{A}}^{-1}\tilde{{\mathbf{\Sigma}}}^{TME}({\mathbf{A}}^{-1})^{T}$, ${\mathbf{A}}_{0}={\mathbf{\Sigma}}^{(out)}_{L_{*},L_{*}^{\perp}}({\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}})^{-1}$ and the definition of $g_{1}$ yields
$g_{1}({{\mathbf{\Sigma}}}^{TME})=g_{1}\left({\mathbf{A}}^{-1}\begin{pmatrix}\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*},L_{*}}&0\\ 0&\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}{\mathbf{A}}^{-1}\right)=g_{1}\begin{pmatrix}\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*},L_{*}}+{\mathbf{A}}_{0}\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{A}}_{0}^{T}&\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{A}}_{0}^{T}\\ {\mathbf{A}}_{0}\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}&\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}=\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*},L_{*}}.$
Combining this formula and the bound $\|{\mathbf{A}}^{-1}\mathbf{x}\|\geq\|\mathbf{x}\|/\|{\mathbf{A}}\|$ results in
$\kappa_{1}=\frac{\sigma_{d}(g_{1}({\mathbf{\Sigma}}^{TME}))}{\sigma_{1}({\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})}\geq\frac{1}{\|{\mathbf{A}}^{-1}\|^{2}\|{\mathbf{A}}\|^{2}}\frac{\sigma_{d}(\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*},L_{*}})}{\sigma_{1}(\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})}.$
Similarly,
$\kappa_{2}=\frac{\sigma_{1}({\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})}{\sigma_{D}({\mathbf{\Sigma}}^{TME})}\leq\|{\mathbf{A}}^{-1}\|^{2}\|{\mathbf{A}}\|^{2}\frac{\sigma_{1}(\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})}{\sigma_{D}(\tilde{{\mathbf{\Sigma}}}^{TME})}=\|{\mathbf{A}}^{-1}\|^{2}\|{\mathbf{A}}\|^{2}\frac{\sigma_{1}(\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})}{\sigma_{d}(\tilde{{\mathbf{\Sigma}}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})}.$ (69)
[…]
It remains to investigate the RHS of (3), that is,
$\frac{d\kappa_{in}{\mathcal{A}}}{n_{1}}\left(\kappa_{in}+\frac{{\mathcal{A}}}{\frac{n_{1}}{d}-\gamma\frac{n_{0}}{D-d}}+\frac{\kappa_{2}{\mathcal{A}}}{\gamma{\mathcal{S}}}(1+\kappa_{in})\right)=\kappa_{in}\frac{d{\mathcal{A}}}{n_{1}}\left(\kappa_{in}+\frac{\frac{d{\mathcal{A}}}{n_{1}}}{1-\gamma\text{DS-SNR}}+\frac{{\mathcal{A}}}{{\mathcal{S}}}\frac{\kappa_{2}}{\gamma}(1+\kappa_{in})\right).$
By the analysis of the Generalized Haystack Model in Section 2, $\frac{{\mathcal{A}}}{{\mathcal{S}}}$ and $\frac{d{\mathcal{A}}}{n_{1}}$ have upper bounds that only depend on $\kappa_{in}$, $\kappa_{out}$, $\mu$ and DS-SNR as follows:
$\frac{{\mathcal{A}}}{{\mathcal{S}}}\leq\frac{C_{\mu_{out},u}}{C_{\mu_{out},l}}\frac{D}{D-d-2}\leq\sqrt{\kappa_{out}}\frac{D}{D-d-2}\leq\frac{\sqrt{\kappa_{out}}}{\mu},$
where the estimation of ${\mathcal{A}}$ follows (9), the estimation of ${\mathcal{S}}$ follows (10), and $C_{\mu_{out},l}$ and $C_{\mu_{out},u}$ are defined in (5). In addition,
$\frac{d{\mathcal{A}}}{n_{1}}\leq\frac{dn_{0}C_{\mu_{out},u}\frac{1}{D-d-2}}{n_{1}}\leq\text{DS-SNR}\frac{C_{\mu_{out},u}(D-d)}{(D-d-2)}\leq\text{DS-SNR}\frac{\kappa_{out}(D-d)}{(D-d-2)}\leq 3\text{DS-SNR}\kappa_{out},$
where the last step follows from the assumption $D-d>2$, so $D-d\geq 3$ and $\frac{D-d}{D-d-2}\leq 3$.
[…]
3.2.1 Proof of Lemma 10
(a) Recall that the TME solution is the minimizer of the objective function
$F({\mathbf{\Sigma}})=\mathbb{E}_{\mathbf{x}}\log(\mathbf{x}^{T}{\mathbf{\Sigma}}^{-1}\mathbf{x})+\frac{1}{D}\log\det({\mathbf{\Sigma}}),$
Recall that using the basis vectors for $L_{*}$ and $L_{*}^{\perp}$, we may write ${\mathbf{\Sigma}}$ as
${\mathbf{\Sigma}}=\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}&{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}\\ {\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}.$
We define
${\mathbf{\Sigma}}^{\star}=\begin{pmatrix}{\mathbf{\Sigma}}_{L_{*},L_{*}}&-{\mathbf{\Sigma}}_{L_{*},L_{*}^{\perp}}\\ -{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}}&{\mathbf{\Sigma}}_{L_{*}^{\perp},L_{*}^{\perp}}\end{pmatrix}$
and note that $F({{\mathbf{\Sigma}}})=F({\mathbf{\Sigma}}^{\star})$ and, by symmetry, for each $\mathbf{x}=[\mathbf{x}_{L_{*}},\mathbf{x}_{L_{*}^{\perp}}]$, there is an equal chance of ${\mathbf{x}^{\star}}=[\mathbf{x}_{L_{*}},-\mathbf{x}_{L_{*}^{\perp}}]$ in the dataset, and $\mathbf{x}^{T}{\mathbf{\Sigma}}^{-1}\mathbf{x}=\mathbf{x}^{\star\,T}({\mathbf{\Sigma}}^{\star})^{-1}\mathbf{x}^{\star}$ and $\mathbf{x}^{T}({\mathbf{\Sigma}}^{\star})^{-1}\mathbf{x}=\mathbf{x}^{\star\,T}{\mathbf{\Sigma}}^{-1}\mathbf{x}^{\star}$.
By the geodesic convexity of $F$ [13, 11], the geometric mean of ${{\mathbf{\Sigma}}}^{(TME)}$ and ${\mathbf{\Sigma}}^{(TME)\,\star}$ should also be the minimizer of $F$. However, for any ${\mathbf{\Sigma}}$, the geodesic mean of ${\mathbf{\Sigma}}$ and ${\mathbf{\Sigma}}^{\star}$, denoted by ${\bf{M}}$, should be “block diagonal” in the sense that ${\bf{M}}_{L_{*},L_{*}^{\perp}}={\mathbf{0}}$ and ${\bf{M}}_{L_{*}^{\perp},L_{*}}={\mathbf{0}}$. One can obtain this property by applying the geometric mean algorithm in [3, (4)], which is block diagonal in every iteration. Since the TME solution is unique (up to scaling), the TME solution is block diagonal.
(b) WLOG we may assume that ${\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}}={\mathbf{I}}$. If not, we may apply $T(\mathbf{x})=P_{L_{*}}\mathbf{x}+{\mathbf{U}}_{L_{*}^{\perp}}({\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}})^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}$ so that this condition is satisfied.
First of all, we will show that ${\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}$ is a scalar multiple of ${\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}}$. Applying the rotational invariance of the dataset within $L_{*}^{\perp}$, whenever ${\mathbf{\Sigma}}^{TME}$ is a minimizer of $F$, for any semi-orthogonal matrix ${\mathbf{U}}\in{\mathbb{R}}^{(D-d)\times(D-d)}$,
$\tilde{{\mathbf{\Sigma}}}^{TME}=\begin{pmatrix}{\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}}&0\\ 0&{\mathbf{U}}^{T}{\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}{\mathbf{U}}\end{pmatrix}$
is also a minimizer. Considering the uniqueness of the solution, ${\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}$ is a scalar multiple of ${\mathbf{I}}={\mathbf{\Sigma}}^{(out)}_{L_{*}^{\perp},L_{*}^{\perp}}$.
Second, we will show that ${\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}}$ has a bounded condition number. We let $\tilde{\mathbf{x}}=({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}{\mathbf{U}}_{L_{*}}^{T}\mathbf{x}$ and note that (20) implies
$\frac{n_{1}}{N}\mathbb{E}_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{T}}{\|\tilde{\mathbf{x}}\|^{2}}+\frac{n_{0}}{N}\mathbb{E}_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{T}}{\|\tilde{\mathbf{x}}\|^{2}+\|({\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}\|^{2}}=\frac{1}{D}{\mathbf{I}}.$ (70)
Since for $\mathbf{x}\in{\mathcal{X}}_{in}$, $\tilde{\mathbf{x}}$ is sampled from a Gaussian distribution with covariance $({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(in)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}$, $\mathbb{E}_{\mathbf{x}\in{\mathcal{X}}_{in}}\frac{\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{T}}{\|\tilde{\mathbf{x}}\|^{2}}$ has the same eigenvectors as $({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(in)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}$, and $\mathbb{E}_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{T}}{\|\tilde{\mathbf{x}}\|^{2}+\|({\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})^{-0.5}{\mathbf{U}}_{L_{*}^{\perp}}^{T}\mathbf{x}\|^{2}}$ has the same eigenvectors as $({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(out)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}$. In addition, by (70), the largest eigenvector of $({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(in)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}$ should also the smallest eigenvector of $({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(out)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}$. Let ${\mathbf{u}}$ be the largest eigenvector of $({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(in)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}$, then its eigenvalue is in the order of $O(\sigma_{d}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-1})$, and
${\mathbf{u}}^{T}\Big{(}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(out)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}\Big{)}{\mathbf{u}}=O(\sigma_{d}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-1}).$
However, ${\mathbf{u}}$ should also be the smallest eigenvector of $({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}[{\mathbf{\Sigma}}^{(out)}]_{L_{*},L_{*}}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-0.5}$, which should be in the order of $O(\sigma_{1}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-1})$, so $O(\sigma_{1}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-1})=O(\sigma_{d}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})^{-1})$ and the condition number of ${\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}}$ is bounded.
Lastly, let us bound $\sigma_{d}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})/\sigma_{1}({\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})$. Assuming that ${\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}={\mathbf{I}}$ and $\sigma_{1}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})=1/\eta_{0}$, then the TME condition implies
$\mathbb{E}_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\mathbf{x}_{L_{*}^{\perp}}\mathbf{x}_{L_{*}^{\perp}}^{T}}{{\|\mathbf{x}_{L_{*}^{\perp}}\|^{2}}+\eta_{0}{\|\mathbf{x}_{L_{*}}\|^{2}}}\succeq\mathbb{E}_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{[\mathbf{x}\mathbf{x}^{T}]_{L_{*}^{\perp},L_{*}^{\perp}}}{\mathbf{x}^{T}({\mathbf{\Sigma}}^{TME})^{-1}\mathbf{x}}=\frac{1}{D}({\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}})=\frac{1}{D}{\mathbf{I}}.$
Taking traces from both sides, we have
$\mathbb{E}_{\mathbf{x}\in{\mathcal{X}}_{out}}\frac{\mathbf{x}_{L_{*}}^{T}\mathbf{x}_{L_{*}^{\perp}}}{{\|\mathbf{x}_{L_{*}}\|^{2}}+\eta_{0}{\|\mathbf{x}_{L_{*}}\|^{2}}}\geq\frac{(D-d)}{D}.$
As a result, $\eta_{0}$ should be chosen such that for $X_{1}\sim\chi^{2}_{D-d}$ and $X_{2}\sim\chi^{2}_{d}$
$\mathbb{E}\frac{X_{1}}{X_{1}+\kappa_{out}\eta_{0}X_{2}}\geq\frac{N(D-d)}{n_{0}D}=1-(1-\text{DS-SNR})\frac{d}{D},$
or equivalently,
$\mathbb{E}\frac{\eta_{0}X_{2}}{\frac{1}{\kappa_{out}}X_{1}+\eta_{0}X_{2}}\leq(1-\text{DS-SNR})\frac{d}{D},$
which implies
$\mathbb{E}\frac{\eta_{0}X_{2}}{X_{1}+\eta_{0}X_{2}}\leq{\kappa_{out}}(1-\text{DS-SNR})\frac{d}{D}.$
Combining it with
$\mathbb{E}\frac{\eta_{0}X_{2}}{X_{1}+\eta_{0}X_{2}}\geq\mathbb{E}_{X_{2}}\frac{\eta_{0}X_{2}}{\mathbb{E}_{X_{1}}X_{1}+\eta_{0}X_{2}}=\mathbb{E}_{X_{2}}\frac{\eta_{0}X_{2}}{D-d+\eta_{0}X_{2}}\geq\frac{\eta_{0}d/2}{(D-d)+\eta_{0}d/2}\Pr(X_{2}>d/2)\geq 0.5\frac{\eta_{0}d/2}{(D-d)+\eta_{0}d/2},$
yields
$\eta_{0}\leq\frac{2\kappa_{out}(1-\text{DS-SNR})d}{D-2\kappa_{out}(1-\text{DS-SNR})d}\frac{2(D-d)}{d}=\frac{4\kappa_{out}(1-\text{DS-SNR})(D-d)}{D-2\kappa_{out}(1-\text{DS-SNR})d}.$
Since ${\mathbf{\Sigma}}^{TME}_{L_{*}^{\perp},L_{*}^{\perp}}={\mathbf{I}}$, $\sigma_{1}({\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}})=1/\eta_{0}$, and ${\mathbf{\Sigma}}^{TME}_{L_{*},L_{*}}$ has a bounded condition number, part (b) is proved. ∎
Acknowledgement. FY and GL were supported by NSF award DMS-2124913 and TZ was supported by NSF award DMS-2318926.